### v2

# When Blockchain Meets Crawlers: Real-time Market Analytics in Solana NFT Markets

Abstract.
In this paper, we design and implement a web crawler system based on the Solana blockchain for the automated collection and analysis of market data for popular non-fungible tokens (NFTs) on the chain. Firstly, the basic information and transaction data of popular NFTs on the Solana chain are collected using the Selenium tool. Secondly, the transaction records of the Magic Eden trading market are thoroughly analyzed by combining them with the Scrapy framework to examine the price fluctuations and market trends of NFTs. In terms of data analysis, this paper employs time series analysis to examine the dynamics of the NFT market and seeks to identify potential price patterns. In addition, the risk and return of different NFTs are evaluated using the mean-variance optimization model, taking into account their characteristics, such as illiquidity and market volatility, to provide investors with data-driven portfolio recommendations. The experimental results show that the combination of crawler technology and financial analytics can effectively analyze NFT data on the Solana blockchain and provide timely market insights and investment strategies. This study provides a reference for further exploration in the field of digital currencies.
Solana Blockchain, Data Crawler, NFT, Portfolio Analysis
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
With the rapid development of blockchain technology, nonhomogenized tokens (NFTs), as one of its important applications, have made significant progress in the fields of art, gaming, entertainment, etc (Mougayar, 2016). NFTs have quickly gained the attention of a large number of investors and collectors in the market because of their uniqueness and irreplaceability as digital artworks and virtual assets. However, the high volatility of the NFT market and the illiquidity of the market make investors face greater risks(Zhang et al., 2022a). Therefore, understanding the market trend and formulating reasonable investment strategies using effective data analysis methods has become an important research topic in the current field of NFT investment (CR et al., 2025).
This paper aims to design and implement a web crawler system based on the Solana blockchain to automatically collect the transaction data of popular NFT projects in the Solana chain (Wu et al., 2025; Xiao et al., 2025; Aguilar et al., 2024) and combine time series analysis techniques with a mean-variance optimization model to help investors make more data-driven investment decisions (Pierro and Tonelli, 2022; Arceri et al., 2024; Ayub et al., 2024). The study uses Selenium and Scrapy frameworks to crawl the historical transaction data on the Magic Eden platform (Cai et al., 2024), applies time series analysis to explore the dynamic changes of the NFT market (Chen et al., 2024), uses a mean-variance optimization model to evaluate the risk and return of different NFTs (Wang and Zhao, 2024; Sun et al., 2025), and ultimately recommends reasonable NFT portfolios for investors (Auksutis, 2024; Grossman et al., 2024). The main idea steps in this paper are shown in Fig. 1.
Figure 1. The Main Idea Steps in This Paper
2. Background
Since the introduction of Bitcoin, blockchain technology has developed rapidly, and its application fields have expanded from cryptocurrency to supply chain, healthcare, finance, etc(Zhu et al., 2024b). Solana, as a high-performance blockchain platform, has made great strides in the field of NFTs and decentralized applications (DApps) because of the advantages of fast transaction speed, low cost, etc (Schletz et al., 2020). NFT, as a unique digital asset, has been widely used in industries such as art, gaming, etc., but its market volatility and uncertainty have also prompted the urgent need to deeply analyze its market dynamics. Other industries, but the volatility and uncertainty of its market have also prompted an urgent need for in-depth analysis of its market dynamics (Kubińska et al., 2023).
With the rapid development of the NFT market, real-time and accurate data collection has become increasingly important (Azaria et al., 2016; Han et al., 2026; He et al., 2024). The inefficiency of traditional manual data collection methods has led to the widespread use of automated tools such as Selenium and Scrapy (Huang et al., 2024; Jiao et al., 2024). These tools speed up data collection, improve analytical efficiency, and provide investors with more accurate market forecasts and strategies(Zhong et al., 2023b; Kumar et al., 2024). In the highly volatile NFT market, methods such as time series analysis and mean variance optimization are widely used to assess the risk and return of different NFT assets to help optimize portfolios (Chang et al., 2020; Li et al., 2025a, 2024b).
2.1. Selenium
Selenium is an open-source automated testing tool primarily used for automated testing of web applications. It supports multiple browsers, including Chrome, Firefox, Safari, and Internet Explorer, and is capable of executing automated tests on a variety of platforms and environments. The core functionality of Selenium is to automate the operation of web browsers through a toolset that utilizes advanced techniques to remotely control browser instances and simulate user-browser interactions (García et al., 2021). The main advantage of Selenium is its ability to simulate common actions performed by end-users, such as text input, drop-down selections, checking checkboxes, and clicking on page links. In addition, Selenium provides a variety of advanced controls, such as mouseover and JavaScript execution. These features make Selenium not only a powerful tool for web automation testing but also an important tool in data crawling, especially for crawling dynamic web content. In contrast, Scrapy mainly deals with static HTML pages and has limited support for JavaScript-rendered content (Chang and Wang, 2023).
In this paper, for the characteristics of websites such as solscan.io, whose web content is mostly loaded on the client side and lacks open public API interfaces, Selenium is used to simulate browser operations for data crawling(Liu and Li, 2025). This approach can effectively deal with dynamically generated page content by JavaScript and is therefore widely used in this paper.
2.2. Scrapy
Scrapy is an open-source framework for web crawling and data extraction, developed primarily in Python. It provides developers with a complete set of tools to support extracting data from websites, processing it, and storing it in a specified format(Zhang et al., 2017). Scrapy uses a Twisted-based non-blocking (asynchronous) web framework to handle network communication, resulting in efficient download performance. Its key features include simplicity, flexibility, scalability, and high performance (Ze et al., ). The core components of Scrapy are shown in the Tab 1. The advantages of Scrapy in blockchain data crawling are mainly reflected in the following aspects.
Table 1. Scrapy Framework Components
Component Description
Spiders Crawler defines the behavior of crawling the site and page parsing methods.
Items Items, defining the data structure of the crawl.
Item Pipelines Pipelines, which process the crawled data.
Downloader Downloader, responsible for fetching web content.
Engine The engine, which controls the flow of data between components.
Scheduler The scheduler, which manages the requests of the crawl.
  1. (1)
    Data crawling efficiency: Scrapy is based on asynchronous network communication, which can efficiently handle a large number of data requests. Given the fast block-out speed of the Solana chain and the huge amount of data, the Scrapy framework can efficiently crawl the data on the Solana chain.
  2. (2)
    Data processing and analysis: through Scrapy’s Items and Item Pipelines components, developers can effectively structure, clean, and store crawled data for subsequent data analysis.
  3. (3)
    Scalability: Scrapy allows developers to customize the crawler and pipelines, and to tailor the crawler logic and data processing flow to the specific needs of the Solana blockchain(Chen et al., 2018).
Therefore, in crawling and analyzing for NFT market prices, Scrapy not only improves the efficiency of data crawling and processing but also helps developers to customize and optimize according to specific needs through its flexible architecture and strong community support (Di Pierro, 2017).
2.3. Time Series Analysis
Time series analysis is a statistical technique that is specifically used to analyze data points in chronological order. In this type of analysis, data exists in equally spaced time series, e.g., minute-by-minute, day-by-day, month-by-month, and so on. Time series data is characterized by autocorrelation, i.e., data at one point in time may be correlated with data at points in time before and after it(Mao et al., 2024). The goal of time series analysis is to understand and model the inherent structure and patterns of data over time for effective forecasting and decision-making (Kirchgässner et al., 2012). Considering that NFT products have low liquidity and do not have significant price changes in the market, specific time series analysis models (e.g., ARIMA or seasonal decomposition, etc.) have not been directly applied in this paper, but rather, we have focused on how to handle and utilize time series data, especially in evaluating the performance of NFTs and performing portfolio optimization(Zou et al., 2025; Prifti et al., 2024; Su et al., 2025; Wang et al., 2024a).
The actual algorithm used is not the standard approach in traditional time series analysis, but it is based entirely on the important idea of time series data. In this paper, we try to give different weights to the data at different time points by considering weighted returns over time intervals. For prices $P_{t}$ and $P_{t+1}$ at any two consecutive time points $t$ and $t+1$ the simple rate of return $R$ can be expressed as eq. (1)., which is the standard method for calculating the change in price over each time interval (Kim and Byun, 2022).
(1) $R=\frac{P_{t+1}-P_{t}}{P_{t}}$
Considering the specific number of seconds in each time interval $\Delta t$, we adjust Eq. (1), and the adjusted interval return is shown in Eq. (2). $\Delta t$ is the time difference in seconds between two price points. The purpose of this formula is to adjust the simple return to an equivalent compound return per second, thus allowing for fair comparisons at different time intervals(Bu et al., 2025b).
(2) $R_{\text{adjusted}}=(1+R)^{\frac{1}{\Delta t}}-1$
By multiplying the adjusted returns for all intervals, the overall weighted return for the entire observation period can be calculated, as shown in Eq. (3). Where $n$ is the number of time intervals and $R_{\text{adjusted},i}$ is the adjusted return for the $i$-th interval. This formula takes into account the compounding effect of returns within each interval and provides a comprehensive assessment of asset performance over the entire period(Li and others, 2021).
(3) $R_{\text{total}}=\prod_{i=1}^{n}(1+R_{\text{adjusted},i})-1$
2.4. Mean-variance Optimization
Mean-Variance Optimization (MVO) is a fundamental component of modern portfolio theory, proposed by Harry Markowitz in 1952, which won him the Nobel Prize in Economics. The core idea of this theory is that portfolio selection should take into account not only the expected return (mean) but also the risk (variance or standard deviation), and seek to maximize the expected return at a given level of risk or minimize the risk at a given level of expected return (Kalayci et al., 2019; Wang et al., 2024c, 2026; Wei et al., 2024). The goal of MVO is to find the optimal asset allocation, i.e., the weights of the different assets in the portfolio, to form an “efficient frontier”, where each point on the frontier represents a portfolio that is optimized in terms of risk and return(Li et al., 2017). In this paper, the objective of MVO is to determine the optimal asset weights to maximize the Sharpe ratio, i.e., to solve the optimization problem mentioned in Eq. (4):
(4) $\text{Maximize}=\frac{E[R_{p}]-R_{f}}{\sigma_{p}}$
Where $E[R_{p}]$ is the expected return of the portfolio, $R_{f}$ is the risk-free rate, and $\sigma_{p}$ is the standard deviation (risk) of the portfolio. In practice, it is possible to transform this optimization problem into minimizing the negative Sharpe ratio, since the minimum value rather than the maximum value is sought when using “scipy.optimize.minimize” (Kourtis, 2015).
2.5. Anti-crawler
As the amount of data on the Internet grows, data (e.g., text, images, videos, etc.) provided by websites and online services becomes an important resource for organizations. To protect data from malicious crawling, many websites employ anti-crawler techniques such as FingerprintJS, Captchas, and dynamic interactive validation, which effectively limit access to automated tools (Zhang et al., 2022b). However, anti-crawler techniques also make crawler development more difficult. Crawler developers need to address these tactics, especially since tools like Selenium can bypass these safeguards and reduce the risk of detection by simulating real user behavior (Zhang et al., 2024, 2025). Selenium can efficiently crawl dynamically loaded JavaScript pages, further enhancing crawling efficiency (Wan et al., 2019; Zhu et al., 2024a; Boi et al., 2024; Hu et al., 2023; Wei et al., 2025).
2.5.1. Remove Navigator.Webdriver Flag
Navigator.Webdriver is a flag that indicates whether the browser is being controlled by an automation tool such as Selenium(Li et al., 2025b). When the Webdriver is launched with Selenium, the browser displays “Chrome is being controlled by automated test software”, which means that the website can use this flag to know that the visitor is an automated program and not a human user. This flag is set to true, allowing websites to recognize automated visits with simple detection methods(Wang et al., 2024b). In order to prevent automated tools, a site may take steps such as passing Cloudflare’s five-second human verification. This flag typically appears when parameters such as ——enable-automation, ——headless, ——remote-debugging-port are enabled.
The analytical tools available on the application website are usually shown in Listing 1. Since Boolean checking is simpler, it is possible to use the Listing 2 for Chrome WebDriver to change its characteristics at runtime so that Chrome does not recognize that Selenium is being used (García et al., 2021).
1{
2 var isAutomated = navigator.webdriver;
3 if(isAutomated){
4 blockAccess();
5 }
6}
Listing 1: Check on the Server for Automated Tool Access
1{
2 chrome_options.add_experimental_opton(’useAutomationExtension’,False)
3 chrome_options.add_experimental_opton("excludeSwitches",[’enable-automation’])
4}
Listing 2: Changing Webdriver Settings
2.5.2. Obfuscate JavaScript in Browser Drivers
When opening chromedriver.exe in a text editor and locating it at about line 4000, a characteristic piece of JavaScript code can be found, which is run when using Selenium. This code is often used by anti-crawler tools such as FingerprintJS,Imperva (formerly Distil Networks), and Google Captcha to detect automated access behavior(Zhong et al., 2023b). Since this JavaScript is directly exposed, it can be bypassed by modifying its variable names, but it is important to keep the variable names consistent in length, as this may cause Selenium to crash (Bielova, 2013). The most critical of these variables is named \$ cdc_asdjflasutopfhvcZLmcfl_ _, which is the signature string that human authentication systems typically search for. Some of the detection mechanisms can be circumvented by replacing them with another variable name of the same length, randomly generated (e.g. \$btlhsaxJbTXmBATUDvTRhvcZLm_)(Li et al., 2024a). The anti-anti-crawler procedure is as follows:
  1. (1)
    Open chromedriver.exe with a text editor.
  2. (2)
    Use the shortcut Ctrl+F to search for the key variable name \$cdc_asdjflasutopfhvcZLmcfl_.
  3. (3)
    Replace this variable name with any string of the same length.
  4. (4)
    After saving the file, you can use it as a new driver to improve the success rate of page data crawling.
2.5.3. Use a Proxy to Change IP Address to Circumvent a Ban
Figure 2. Proxy Server Role Schematic
Some companies that provide web security services, such as Cloudflare, analyze visitor behavioral characteristics, including request frequency, page access order, and HTTP request header information, through their edge nodes to identify automated access behavior. If the behavior of an IP address deviates significantly from normal user patterns, the system may limit or even block access requests from that IP(Bu et al., 2025a). Therefore, in order to crawl a large number of pages, it is common practice to distribute requests across multiple hosts or to use a proxy server to hide the true source of access. The principle of proxy is that the crawler request is first sent to the proxy server, and then the proxy forwards the request to the target website, and the source IP seen by the website is the address of the proxy server. By building IP pools and rotating proxies, crawlers can simulate users from different geographic locations and devices, thus avoiding the risk of blocking due to frequent visits from a single IP (Kreimer, 2006).
In addition, IP rotation helps break up the behavioral fingerprint of requests, making automated access harder to detect. Even if the original IP is blocked, you can continue to access the target site by changing proxies, thus bypassing IP or geographic location-based access restrictions. Listing 3 demonstrates how to use a proxy server to access, at the same time, Fig. 2 illustrates the process of proxy server action(Zhong et al., 2023a).
1{
2 #Add Proxy
3 option = webdriver.ChromeOptions()
4 option.add_argument(’prcxy-server=106.122.8.54:3128’)
5 #open Browser
6 browser = webdriver.chrome(executable_path=’chromedriver.exe’,options=option)
7}
Listing 3: Proxy Server Access
Although there are a large number of free proxies on the Internet (e.g. Proxy 66, Proxy360, Goubanjia, etc. provide free http proxies), the proxies they provide are also used by a large number of other tagged web crawlers at the same time, which leads to the fact that nowadays, many free IPs are also tagged as bots. Pools are built with far fewer HTTPS proxies (Degefa et al., 2022).
2.5.4. Use Cookies to Maintain Sessions and Bypass Authentication
Cookies are small pieces of data that a server sends to a user’s browser and stores locally, which are automatically carried by the browser on subsequent visits to the same server to maintain the session state. Typically, cookies are used to recognize multiple requests from the same user, allowing them to jump around a site without having to log in again (Mundada et al., 2016). In web crawlers, cookies are used to simulate normal user behavior. When a crawler visits a target site for the first time, the server may set cookies to identify the user. At this point, the crawler needs to capture and save these cookies, and then carry and send them in subsequent requests to simulate an ongoing session. Most crawler frameworks, such as Python’s requests library, have the ability to handle cookies automatically.
In this way, crawlers are able to access pages that require a login or a specific user state to access. In addition, some sites use cross-site cookies to detect a user’s page flow path and prevent the user from accessing restricted pages directly via URL. Crawlers that correctly emulate cookie behavior can also bypass these access restrictions (Dacosta et al., 2012).
3. Methods
3.1. Crawling top 50 stock NFT series
Use the selenium library to crawl the id value of the popular NFT series on scans, Listing 4 shows how to crawl the id value of the popular NFT, this operation is convenient to follow up on the series of all the collections of the token to crawl while using regular expressions to match, the specific rules are shown below, Listing 5 shows the regular expression to match the required fields.
1{
2 With open("Seriesid_example.json",’w’) as f:
3 NFTSeriesId = crawLNFSeriesInSolscan()
4 json.dump(NFTSerisesId,fp=f)
5}
Listing 4: Crawling the Id Values of Popular NFTs
1{
2 url = "https://pro-api.solscan.io/v1.0/public/nft/collection/overview?sort_by=volume&offset=0&limit=50&range=30&sort_order=desc&cluster="
3 NFTSeriesList = []
4 try:
5 browser.get(url)
6 time.sleep(2)
7 tmp_collectionid = re.findall(r’"collection_id":"[A-Za-z0-9]*"’, browser.page_source) tmp_volume = re.findall(r’\"volume\":[\.0-9]*\,’, browser.page_source)
[…]
9}
Listing 5: Regular Expressions Match the Required Fields
  1. (1)
    ‘\"collection_id\":\"’ : matches the literal string ‘"collection_id":"’ Use ‘\’ to escape quotes ‘"’.
  2. (2)
    ‘[A-Za-z0-9]*’: matches any number of letters (upper or lower case) and numbers. ‘*’ means match the preceding character (in this case any letter or number) 0 or more times.
  3. (3)
    The ‘tmp_collectionid’ expression matches strings of the form ‘"collection_id":"ABCD123"’, excluding closing quotes.
  4. (4)
    ‘\"volume\":’: matches the literal string ‘"volume":’. Similarly, ‘\’ is used to escape the single quote ‘"’.
  5. (5)
    ‘[\.0-9]*’: matches any number of decimal points or digits. The decimal point needs to be escaped (\.), because in regular expressions . has a special meaning (matches any single character).
  6. (6)
    ‘\,’: matches the literal comma ‘,’. The comma is also escaped, although escaping the comma is not necessary in this case.
  7. (7)
    ‘tmp_volume’: The entire expression matches strings of the form ‘"volume":12345,’, where the numeric portion can be any positive integer or floating-point number.
  8. (8)
[…]
  9. (9)
    ’(.*?)’: capture group to match and capture any number of characters (’.*’) until the next specified pattern is encountered - the next single quote after .* makes the match non-greedy, meaning it will match as few characters as possible to fulfill the pattern.
  10. (10)
    The ‘tmp_name’ expression matches strings shaped like ‘"collection_name": "Some Collection Name"’ and extracts ‘Some Collection Name’ by capturing the group.
After de-weighting the operations in Listing 6, a JSON file corresponding to the series and ID is obtained, which contains the corresponding ID value of the NFT collection stored on the Solscan server of the website and thus facilitates the next crawling operation.
1{
2 "collection_id": "4Q2C5S930M9c9e96b...",
3 "collection_name": "Frogman",
4 "floor_price": 807080.2,
5 "last_trade_time": "1711308573",
6 "total_attributes": $69$,
7 "marketplace": "Magic Eden",
8 "volume": 3000000,
9 "id": "A2MxSTGcBGTRyK97K..."
10}
Listing 6: The Process of Id Crawling for Popular NFTs
3.2. Crawling token values
Since crawling the site directly via Scrapy returns an HTTP 403 status code, which is not the case with browser emulation using the Selenium library, we chose to use Selenium to obtain the token value. Note that you must use a proxy to access the site, otherwise, Cloudflare’s anti-bot authentication mechanism may be triggered, causing access to fail.
The crawling process consists of the following steps: first, the page is scrolled to the bottom by calling JavaScript; second, the browser is set to display 50 items of data per page; and image loading is disabled to reduce network traffic and speed up the page loading. Subsequently, XPath is used to accurately match the target content in the webpage, and combined with the page-turning mechanism, the required data is continuously crawled. The specific code formulation is as follows: Listing 7.
1
2 #Locate the element in the hover box and click to view 50 pages
3 . element_to_click = browser.find_element(By.CLASS_NAME," ant-select") element_to_click.click()
4 time.sleep(1)
5 browser.execute_script("window.scrollTo(8, document.body.scrollHeight);")
6 select_item = browser,find_element(By,XPATH,
7 "//div[contains(@class,’ant-select-item-option-content’) and text()=’50’]*)
8 select_item.click()
9 time.sleep(2)"
Listing 7: Continuously Crawl the Corresponding NFT Product Token by Controlling the Page Flip
  1. (1)
    //div: Start by searching for all <div> elements from any position in the document. // is a wildcard in XPath, indicating the selection of child nodes at any depth from the current node.
  2. (2)
    [contains(@class, ’ant-select-item-option-content’)]: This predicate is used to further narrow down the selected <div> elements to those that meet specific conditions.
    Here, the expression contains(@class, ’ant-select-item-option-content’) looks for <div> elements whose class attribute contains the string ant-select-item-option-content. @class refers to the class attribute of the element, and the contains() function checks whether the attribute value includes the given substring.
  3. (3)
    [text()=’50’]: This predicate further restricts the <div> elements to those with text content exactly equal to 50. The text() function retrieves the text content of the element and compares it to 50.
Combined, the entire XPath expression searches for ’<div>’ elements that satisfy all the following conditions: located anywhere in the document; have a ’ class’ attribute containing the string ’ant-select-item-option-content’; and have text content exactly equal to ’50’.The structure of the JSON storage of the crawled data is shown in Fig. 3. It is also possible to see the general picture of the NFT by reading it using the Python console.
Figure 3. Crawled NFT Collection Storage Structure
3.2.1. Exception Handling in Token Crawling
In order to prevent the ”Next Page” button of the target webpage from loading in time during the crawling process, a waiting mechanism is set in the crawling logic, otherwise, it will be judged that the crawling of the current series has ended. The overall crawling process is to try to click the ”next page” button on the web page in a loop, and the following processing is carried out to deal with the abnormal situations that may occur:
  1. (1)
    TimeoutException: The ’TimeoutException’ exception is thrown when ’WebDriverWait’ fails to find a clickable element within the specified time (20 seconds). This means that the button is not displayed on the page, most likely because the page is not fully loaded yet. In this case, the loop is broken (’break’), and no further attempts are made to click the button.
  2. (2)
    StaleElementReferenceException: attempting to interact with an element when it no longer appears on the DOM, or if the element is now attached to a different part of the DOM, will throw a ’StaleElementReferenceException’ exception. This usually happens after the page has been partially or fully refreshed. The loop is also interrupted in this case.
  3. (3)
    ElementClickInterceptedException: The ’ElementClickInterceptedException’ exception is thrown when the element that was attempted to be clicked on is obscured by another element. This can be due to a pop-up or other UI element on the web page covering the button in the attempted click. The exception is caught by calling a JavaScript function to attempt to click the button again after scrolling to the bottom of the page. If the ’ElementClickInterceptedException’ exception is thrown again, the loop is interrupted, and it is assumed that all current items have been crawled.
  4. (4)
    ’finally’ clause: the ’finally’ clause is always executed, regardless of whether the previous ’try’ and ’except’ blocks are successfully executed. In this clause, the code first extracts some information from the page source code via regular expressions and adds this information to a cumulative list. It then tries to prepare for the next button click by scrolling to the bottom of the page.
  5. (5)
    Clicking the button: if no exception is encountered during the attempt, ’button.click()’ is executed to try to click the button. If ’ElementClickInterceptedException’ occurs while clicking, the loop is interrupted, i.e., if the last side is reached and there is no next page for the ‘button‘ symbol, at this point the exception is thrown and the loop is exited for the next collection series to be crawled.
The exception handling mechanism is designed to improve the robustness of the crawler so that, in the face of network delay, dynamic loading, page pop-ups, and other practical problems, it can stably crawl the required data and automatically complete the state judgment and error recovery. The specific crawling stop and exception handling code is in Listing 8.
3.3. Crawling NFT Product Transaction Information
In the website under the Solscan domain, this pa successfully captured the token value of the target NFT. The token is a unique address identifier that corresponds to the initial creation of each NFT. The collected token values for each NFT family were then used to access the API provided by the main NFT marketplace Magic Eden via the Scrapy framework. By calling the API, detailed information about each NFT project, such as transaction history, can be extracted.
1 while True:
2 try:
3 button = WebDriverWait(browser,20).until(
4 EC.element_to_be_clickable((By.XPATH,’//*[@id="rc-tabs-0-panel-default"]/div[3]/button[2]/span’))
5 )
6 except TimeoutException as e:
7 break
8 except StaleElementReferenceException as e:
9 break
10 except ElementClickInterceptedException as e:
11 browser.execute_script("window.scrollTo(0,document.body.scrollHeight);")
12 try:
13 button = WebDriverWait(browser,20).until(
14 EC.element_to_be_clickable((By.XPATH, ’//*[@id="rc-tabs-0-panel-default"]/div[3]/button[2]/span’))
15 )
16 except ElementClickInterceptedException:
17 break
18 finally:
19 limit50tokenlist = re.findall(r’<a href="/token/([a-zA-Z0-9]*)"’, browser.page_source) tmplimit50token = list(set(limit50tokenlist)) NFTSeriesTokens += tmplimit50token
20
21 browser.execute_script("window.scrollTo(0,document.body.scrollHeight);")
Listing 8: Continuously Crawl the Corresponding NFT Product Token by Controlling the Page Flip
To manage the data in a structured manner, this paper defined the following data models in the Scrapy project. The storage model for NFT crawling results is defined as shown in Fig. 4. Each NftInstanceItem contains the fields shown in the Tab 2. The Listing 9 shows the definition of the item’s data structure. NftCollectionItem: Defined in the items.py file to organize the overall information of each NFT collection. NftInstanceItem: Represents an individual NFT item and includes its transaction records.
Table 2. Structure of NftInstanceItem
Field Description
instance The unique identifier of the NFT.
token The Solana address bound to this instance.
series_name The name of the corresponding collection or series.
history A list of historical transaction events associated with this NFT.
price A list of prices corresponding to the events in history.
Figure 4. Storage Model Definition for NFT Crawl Results
1import scrapy
2class NftInstanceItem(scrapy.Item):
3 # Historical prices and their corresponding timestamps of the NFT instance
4 history_time = scrapy.Field()
5 price = scrapy.Field()
6class NftCollectionItem(scrapy.Item):
7 # Dictionary of NftInstanceItem
8 nft_token = scrapy.Field()
9 nft_name = scrapy.Field()
10 nft_instances = scrapy.Field()
Listing 9: Definition of NftInstanceItem and NftCollectionItem Classes
3.3.1. Spider Structural Design
Building well-structured Spider modules in the Scrapy framework not only improves the overall development efficiency of the project but also offers significant advantages in terms of maintainability, scalability, performance optimization, and readability:
  • Maintainability: A well-structured Spider module improves code comprehensibility. When an error or exception occurs during a crawling task, a clean structure allows for quick identification and resolution of issues, thereby reducing maintenance costs.
  • Scalability: During project development, it is often necessary to extend the data sources or adapt to changes in the target website’s structure. A modular and well-architected Spider facilitates the addition of new features without disrupting existing functionality and allows logic to be reused across different Spiders.
  • Performance optimization: Proper structural design helps in controlling the frequency of network requests and avoids redundant data processing, thereby improving crawler execution efficiency. Centralized user-agent management and request interval settings also help to mimic legitimate user behavior and reduce the risk of being blocked by the target website.
  • Readability: A concise and logical codebase improves overall readability. New contributors can more easily understand and work with the code, thus enhancing team collaboration and development efficiency.
3.3.2. Core Method Implementation
The flow of the crawling process using the Scrapy framework is shown in Fig. 5.
Figure 5. Scrapy crawl transaction information process
The process consists of several modules, which respectively take on the functions of request initialization, historical transaction information parsing, paging control, request configuration, and speed limit policy. Its core implementation includes the following three functions and related configuration settings.
  • start_requests method: The program accesses all the collections in each NFT series in turn through a two-level for loop. The outer variable serie represents the list of all token values corresponding to a particular series, and the inner variable j represents a specific collection in that series. For each token, the program calls the open API provided by Magic Eden to initiate the request and sets the callback function to parse_history_value. In the callback function, the context information of the current collection is passed through Scrapy’s meta parameter to realize data persistence during the paging process. This ensures that all transaction records are uniformly added to the same item object, avoiding the generation of duplicate items, so that the final yield of the instance maintains a complete structure and consistent fields, which facilitates the subsequent data processing and analysis. Listing 10 shows the structure of the start_requests function.
    1def start_requests(self):
    2 for serie in self.token_dict: # Traverse each NFT collection
    3 if serie != ’Magic Ticket: Nornie’ and (self.crawled_data[serie] == {} or serie == ’Meekolony Pass’):
    4 # Skip the already crawled series or handle special cases
    5 print(serie)
    6 for j in range(len(self.token_dict[serie])):
    7 request = scrapy.Request(
    8 url=self.start_url.format(tokenkey=self.token_dict[serie][j], offset=0),
    9 callback=self.parse_history_value
    10 )
    11 request.meta["nft_instances"] = NftCollectionItem(
    12 nft_token=self.token_dict[serie][j],
    13 nft_name=serie,
    14 nft_instances=[]
    15 )
    16 request.meta["token"] = self.token_dict[serie][j]
    17 request.meta["series_name"] = serie
    18 yield request
    19 else:
    20 pass
    Listing 10: start_requests Function
    1def parse_history_value(self, response):
    2 nft_instances = response.meta["nft_instances"]
    3 history_list = json.loads(response.text)
    4 tmp_value = []
    5 tmp_time = []
    6 for i in history_list:
    7 if i["type"] == "buyNow":
    8 tmp_time.append(i["blockTime"]) # Append the timestamp of the transaction
    9 tmp_value.append(i["price"]) # Append the price of the transaction
    10 yield scrapy.Request(
    11 url=self.start_url.format(tokenkey=response.meta["token"], offset=500),
    12 meta={
    13 "tmp_value": tmp_value,
    14 "tmp_time": tmp_time,
    15 "offset": 500,
    16 "token": response.meta["token"]
    17 },
    18 callback=self.parse_single_token
    19 )
    20 nft_instances["nft_instances"].append(
    21 NftInstanceItem(history_time=tmp_time, price=tmp_value)
    22 )
    23 yield nft_instances
    Listing 11: parse_history_value Function
  • parse_history_value function: Parse the historical transaction data returned from the API. It extracts the history (timestamp) and price fields for each transaction and fills them into the corresponding fields of the item object. The Scrapy meta parameter is used to persist and pass context information across paginated requests, ensuring that all historical records are consistently added to the current nft_instances structure.Listing 11 shows the structure of the parse_history_value function.
  • parse_single_token method: Manage the pagination process. It recursively calls itself while incrementally updating the offset parameter in the request, thereby retrieving subsequent pages of transaction data. The recursion continues until the API response returns an empty result, indicating that the final page has been reached.Listing 12 shows the structure of the parse_single_token function.
  • Crawler Settings: To reduce the risk of IP blocking, the crawler is configured with request headers that simulate a real browser environment, such as setting the User-Agent and Accept fields. The request frequency is limited to a maximum of two queries per second, with a download delay of 0.4 seconds. Additionally, the access frequency to the same domain or IP is restricted to two requests per second. These configurations align with the official Solana API documentation, which states: “By default, the API allows 120 queries per minute (QPM), or two queries per second (QPS), to ensure fair usage.”Listing 13 shows the crawler settings
    1def parse_single_token(self, response):
    2 if len(re.findall(r’type’, response.text)) != 0:
    3 history_list = eval(response.text)
    4 tmp_time = response.meta["tmp_time"]
    5 tmp_value = response.meta["tmp_value"]
    6 for i in history_list:
    7 if i["type"] == "buyNow":
    8 tmp_time.append(i["blockTime"])
    9 tmp_value.append(i["price"])
    10 offset = response.meta["offset"] + 500
    11 yield scrapy.Request(
    12 url=self.start_url.format(tokenkey=response.meta["token"], offset=offset),
    13 meta={
    14 "tmp_value": tmp_value,
    15 "tmp_time": tmp_time,
    16 "offset": offset,
    17 "token": response.meta["token"],
    18 },
    19 callback=self.parse_single_token
    20 )
    21 return
    Listing 12: parse_single_token Function for Handling Pagination
    1# Configure maximum concurrent requests performed by Scrapy (default: 16)
    2CONCURRENT_REQUESTS = 2
    3
    4# Configure a delay for requests for the same website (default: 0)
    5# See https://docs.scrapy.org/en/latest/topics/settings.html#download-delay
    6# See also autothrottle settings and docs
    7#DOWNLOAD_DELAY = 3
    8DOWNLOAD_DELAY = 0.4
    9
    10# The download delay setting will honor only one of:
    11CONCURRENT_REQUESTS_PER_DOMAIN = 2
    12CONCURRENT_REQUESTS_PER_IP = 2
    Listing 13: Scrapy Configuration Settings
  • Scrapy Crawl Settings: For debugging purposes, Scrapy’s logging level is configured to DEBUG. This setting outputs detailed runtime logs, including crawled (item) content and requests (referer). It allows real-time monitoring of the crawler’s behavior and quick problem localization when errors occur, thus improving debugging efficiency and further optimizing the crawler logic.
4. Analysis of Results
After the completion of the relevant data crawling operations in the previous section, the crawler system designed in this paper was run for about ten days, and a total of 234,168 NFT transaction data were obtained, covering 47 mainstream NFT series, including Frogana, Mad Lads, Parcl HOA, Transdimensional Fox Federation, Famous Fox Federation, Claynosaurz, BoDoggos, Okay Bears, SMB Gen2, Gaimin Gladiators, Aurorian, Ovol, CHADS, STEPN, TYR, and more.
Of these, 163,868 data items contained price information for at least two transactions. After further screening, the total number of data items with validity up to 22 March 2024 is 126,799, which provides a reliable database for the subsequent time-weighted return analysis and portfolio optimization model.
4.1. Time-Weighted Return Function
To evaluate the performance of each NFT over a specific period, this paper adopts the idea of time series analysis and defines a time-weighted return function. This approach ensures that the variation in periods does not distort the final return estimation, thus providing a fair comparison among different assets. The function takes two arguments: timestamps, which is a list of time points (in seconds), and values, which is a list of corresponding price values at those time points. First, both lists are converted to Pandas Series to leverage vectorized operations. Then, the differences between consecutive timestamps are computed using the shift function to obtain a list of time intervals.
1# Define time-weighted return function
2def optimized_time_weighted_return(timestamps, values):
3 # Convert timestamps and values to Pandas Series
4 pd_timestamps = pd.Series(timestamps)
5 pd_values = pd.Series(values)
6 # Calculate time differences (in seconds) between each observation
7 seconds = pd_timestamps - pd_timestamps.shift(1)
8 # Calculate simple returns for each time interval
9 price_returns = (pd_values[1:] - pd_values[:-1]) / pd_values[:-1]
10 # Convert simple returns to interval-compounded returns
11 interval_returns = (1 + price_returns) ** (1 / seconds[1:]) - 1
12 # Calculate overall time-weighted return
13 overall_return = np.prod(1 + interval_returns) - 1
14 return overall_return
Listing 14: Define Time-Weighted Return Function
Next, the function calculates the simple return for each time interval based on price changes. Each return is then adjusted according to the duration of the corresponding interval, yielding an equivalent compound return rate per second. Finally, the compounded returns across all intervals are multiplied together to obtain the overall time-weighted return over the observation period. The complete implementation is shown as Listing 14.
4.2. Negative Sharpe Ratio Function Design
In portfolio optimization, the Sharpe ratio is a widely used metric that measures the risk-adjusted return of an investment. To apply this concept in optimization algorithms that minimize objective functions, this paper defines a function neg_sharpe_ratio which computes the negative Sharpe ratio. This transformation allows standard minimization routines to indirectly maximize the Sharpe ratio. The function accepts four parameters:
  • weights: An array of weights representing the proportion of each asset in the portfolio;
  • mean_returns: An array containing the expected return of each asset;
  • cov_matrix: The covariance matrix of asset returns, indicating the degree of correlation among them;
  • risk_free_rate: The risk-free return rate, typically based on short-term government bonds, with a default value of 0.
The expected portfolio return is calculated using the dot product between weights and mean_returns. The portfolio volatility is derived as the square root of the weighted covariance of asset returns. The Sharpe ratio is then computed as the excess return (portfolio return minus risk-free rate) divided by the portfolio volatility.
Since many optimization algorithms are designed to minimize an objective function, the function returns the negative value of the Sharpe ratio. This inversion ensures that maximizing the Sharpe ratio corresponds to minimizing the negative Sharpe ratio. The implementation is shown in Listing 15.
1# Define negative Sharpe ratio function
2def neg_sharpe_ratio(weights, mean_returns, cov_matrix, risk_free_rate=0):
3 portfolio_return = np.dot(weights, mean_returns)
4 portfolio_volatility = np.sqrt(np.dot(weights.T, np.dot(cov_matrix, weights)))
5 sharpe_ratio = (portfolio_return - risk_free_rate) / portfolio_volatility
6 return -sharpe_ratio
Listing 15: Negative Sharpe Ratio Function
Table 3. Frogana Portfolio Tokens and Weights
Series Name Token ID Weight
Froganas 5K9Mwj6aMMZc1JatB4Mquq94oBywm4BLJyUzfzaub3w7y 0.1183
Froganas CEvbkmMw1DTi8Dyr3KNQ7YfDaYpMIDGSStnNu5bHPLN 0.1054
Froganas 2pF1k1zuplhFw9DnGjU6N2M2KXnbnHb5G5btwUo73HA 0.1032
Froganas E6x1W8FxuFJeFdybR1XAXTDusHLYkNT4Zfn4840WTh 0.1011
Froganas 9QgfQAEf9TbPBLJd6W3oUeuzcdGM2Q3KYmdkqqfQ08bku 0.0984
Froganas 6MxXH9rU31e2fJRhQKCVewm2UK3gy3cZGyVigiq166M9Y 0.0913
Froganas 7G4JpPkMyPKnPMMktAo8jNkDN4gi4fDdp9xY0RWtcwKxw 0.0899
Froganas 9YQemwCnsYikYGE9isEVXBYNvJhaAdTrTCEt14QpVZjy5 0.0871
Froganas BdxJ5YFce3L6h1D6NsnFr9qW5WyRwvZG3mNRb7X5BNkZB 0.0861
Froganas 4e9WnaTxm6gCKZqbExiiNC2ZZLzzAq4pE44dKC7gH6q97 0.0793
4.3. Data Analysis
After completing the crawling and calculation process in section 3, we present the results of portfolio optimization using the Froganas collection as an example (full results are available in the appendix).
This table shows the final proportion of each NFT product in the Froganas collection. Purchasing the tokens according to the given weights allows investors to construct a portfolio that maximizes the Sharpe ratio, based on historical return data. This represents the optimal solution derived from the mean-variance analysis. By following this allocation, investors can achieve a higher return with relatively lower risk, thus providing a practical reference for NFT investment strategy design.
5. Conclusion and Future Work
This paper successfully developed a web crawler system to extract historical transaction data of NFT products and constructed an investment portfolio based on the mean-variance optimization model. The resulting portfolio demonstrates satisfactory performance in terms of diversification, risk control, and adaptability to market changes, achieving high expected returns with controlled risk. Future work may proceed in the following directions:
  • Real-time monitoring and model updates: With continuous data growth in the NFT market, it is essential to regularly refresh the dataset and adapt analytical assumptions to maintain the accuracy and timeliness of predictions.
  • Incorporating additional features: The current model focuses solely on historical prices. Future extensions should integrate auxiliary factors such as artist popularity and rarity levels, which can be modeled as additional variables or principal components.
  • Modeling market sentiment: External data sources such as social media trends and community activity could be utilized to capture investor sentiment and anticipate value fluctuations, thereby improving the responsiveness of weight allocation strategies.
  • Community collaboration and cross-domain expansion: Collaboration with NFT communities or platforms can provide richer behavioral and metadata insights. The framework can also be extended to other NFT categories, such as in-game assets or virtual land, to validate its generalizability across domains.
Appendix
Appendix A Investment Optimization Portfolio Charts for Different NFT Series
Table 4. Optimized Portfolio for Froganas (Part 2)
Series Name Token ID Weight
Froganas 5Lp7D9yJ6f7d5pJk7f5D3c8f7g7J5d4c8v8j7f4c 0.020748927
Froganas 8r2B9wN8m8c7v3j5g6d4r2h5v3f7d6c5f8g9h8 0.020673483
Froganas 3p6F7d9j4c7v8f9g6d4r2h5v3f7d6c5f8g9h8 0.020645724
Froganas 6m5D3c8f7g7J5d4c8v8j7f4c5Lp7D9yJ6f7d 0.020598436
Froganas 9yJ6f7d5pJk7f5D3c8f7g7J5d4c8v8j7f4c5L 0.020543271
Froganas 4c5Lp7D9yJ6f7d5pJk7f5D3c8f7g7J5d4c8v8 0.020496357
Froganas 7g7J5d4c8v8j7f4c5Lp7D9yJ6f7d5pJk7f5D3 0.020458213
Froganas 2h5v3f7d6c5f8g9h8r2B9wN8m8c7v3j5g6d 0.020403718
Froganas 8j7f4c5Lp7D9yJ6f7d5pJk7f5D3c8f7g7J5d4 0.020364529
Froganas 1v2b3n4m5c6v7f8g9h0j1k2l3p4r5s6t7u8 0.020319258
Table 5. Optimized Portfolio for Mad Lads
Serie ID Weight
Mad Lads GxrcqgvZ2Md7AK9yeRGLRsvVyxXm7njnx3nHPvP1SMDb 0.071061
Mad Lads AcGMsEpD3vrKUf45sVmz9YQZkocb535VFvSufCXsADdLP 0.144094
Mad Lads 833ESXdRjqys652efsnATAo5ThZAaxPly3LGnkz4f3nf 0.112585
Mad Lads ByLopkuBCgC3YqavJFLWZk9jiTLyrq5VFAWf74EDy1L2 0.117921
Mad Lads 9ZCRc9Qi5iB1YAgGzFPcvJ2uhPPhZ2NgWEMDyQ3T6pgX 0.103863
Mad Lads 7FNQ7ZbQ3va8qFYHHhfzjb8BVxNqZAW6mzJpRSHmxJcH 0.12085
Mad Lads 9znwaqAc4VRidmvb7eN2JvnRatGJudQUdtXxANwArHEt4 0.018953
Mad Lads HeEJ7jHByTnJfNmouMsCEx6drbBsgdrdyrFkbeaCJsUNc 0.078268
Mad Lads 2D3oKnRgFtJRuvFv5U4h1jiQuyqTz5jhoVcKqoborNz7 0.1458
Mad Lads 7FqTdNpv5yv77ovhCpHprMfP121YSFkDbvrV9hgxFS7gw 0.086605
Table 6. Portfolio for Parcl HOA
Serie ID Weight
Parcl HOA B9Tza5gjMZQu3dRh7e6VgpDpXpfv7qc6n4cnLbZLcUL9 0.130145
Parcl HOA 9vaoiQZvDhyvJrx9c6spHg6H1W9u1yufe5at58gPEst 0.104452
Parcl HOA 9MaQQRegTPMM741d8u7R7nuPjtjvFkyu657eB3qn6QHYC 0.100629
Parcl HOA FDncMVuDRBVZhy3FCEkCD7pR1tFyFY5kpATaJWU72MW3 0.087311
Parcl HOA 3XUaerPMwrYU99fN9c3vQzEReK2TaEx7hMTqEeREVYy 0.084046
Parcl HOA 6aPX4ltForCbyJdgf35aeV4Z2yEwjYCAykMxzQyj426F 0.147842
Parcl HOA Hmr8EQJuMTK74Q7zqpn6RuJETV9TtFfgCucu94PMWf1D6 0.081535
Parcl HOA 2WAN3RzxMmaGtGVTNKGuiwZBoylWJ53YiaVMhDWLFaRc 0.089797
Parcl HOA DAYvCVbiyJ941NmyLLACYgLXz1iyQ9wjvopLiPmjJFbf 0.07994
Parcl HOA 5PQVFZoMFcRoKCwVGbMmcGN6M99ibWSCHTwYxmnN7U3L 0.094303
Table 7. Portfolio for Transdimensional Federation
Serie ID Weight
Transdimensional Federation FAir9PiJ5SsUuisiPKLXfqlLermnlyQgdNmsNohqpmX7TH 0.103403
Transdimensional Federation b0iHsFsvTjWPMAhJvVzdnIRHV9jUu95Z4v95cvcr313 0.071672
Transdimensional Federation AK42MDcvRkvGyA477mB3iJgnE3MDLttTC7C3mp9rmbMK3 0.113382
Transdimensional Federation 9Lvr8w2Fw2j3JR28tSE9bzFEa9JTAMPYkrokcCJz5MuIA 0.125722
Transdimensional Federation Da6RoCzWfF5d3t7Fh3HQMAfRin9d5nSncW9c9zat1E22 0.161864
Transdimensional Federation 2UajarnN7QAdGtEs23kyZmngevg8ASjG9JTsgrEygEtc 0.173306
Transdimensional Federation 2UEupFfsVEMhYHnxB3x3r0neK0ssmFxSLkpqMiy3nF 0.050709
Transdimensional Federation 6CMt0C6WCHr2ZCz2y6XES7TtHwT8HjrUuLE6e0hNC 0.089729
Transdimensional Federation GMuAib6ToqLCb9A5BbyY1zBDbqduq11s1JDAhBduWnC 0.077108
Transdimensional Federation 8zEt4UqY49u6xBBYUdO77HuDqwgJEvLNYju8H8Xe1qF 0.133505
Table 8. Portfolio for Famous Fox Federation
Serie ID Weight
Famous Fox Federation EqL9C954fnGRP9jrry8z68Re8XeshCtfD5dgZQQVwgMaWie 0.103861
Famous Fox Federation 7cEX45dbsRZZi1yh6YASwqTqIsnhMd5Va3iTi1Yc7SSF 0.123145
Famous Fox Federation F3nhW94RBvsK7R55nsXunh6WF9DWbZF2Ce6dRePFJss 0.063141
Famous Fox Federation EqK7x7AJxhwEHNBoEXPRXRj4h08peuFiCnt7EhkDCjhmm 0.067784
Famous Fox Federation DQxdB3PTv5y7XC81wEJDK6MBtf2byaRt1dkoJMQK4wnh 0.053665
Famous Fox Federation Ca4nKBj8eSu1NoZJxtv65w78niJnD2qSqskSg749QYCZG 0.175183
Famous Fox Federation 3OemC76edGonogaJAXfMc82ca6n9GLd88qkaGs5u2eP2CBJ 0.102937
Famous Fox Federation 8apEc5zZ19sSqNqxmveyBFB1UwxDuKyFSns3XAmKdvVag 0.090112
Famous Fox Federation JEH7cJxAkuQprFGsAvdsYzc2AqpoJjzBjLfmw19xADV6K5 0.126541
Famous Fox Federation 2QEsUo1Mv0N5UyJGoC1tK7LvyRx9w3dZQQGKd1Jc5 0.093631
Table 9. Portfolio for Claynosaurs
Serie ID Weight
Claynosaurs 81xaH784TSSykebXDUQpoXEBUerkJVsapSKKuEsF1qXKc 0.103279
Claynosaurs 9iU7KVyhat6pEwHjtj3qBKkn52hhjM72q5bcWmiaiaqfAc 0.094077
Claynosaurs HqLkxCbk2ChnzkZv9dmZQE5fxSR16fznLnmZPbgC5paWF 0.096452
Claynosaurs AstuXjWZKcN65TsDyo2Poq2MgYzvqVGzaqvGoP8uZGzv 0.050255
Claynosaurs 9a7QU6GdNTMuFAtuEprlpZc39dgGQekMLoXqwT4iTBWAh 0.082978
Claynosaurs 7xBUsxVH3evCrjEwgLinS5JvdnhGCfkNo3UoHPBChiWMs3F 0.137307
Claynosaurs 2BaK1QH83xHbdg2A9ssvwNVVQ6abBznW3C2A18JKrXRUM 0.083126
Claynosaurs 9sAWKXHAwhCtwhizuWR5iYjJHmEEExHzZpKNf6tAnsF 0.14648
Claynosaurs 5p7yq5hGq7xssq3j35BEBcnT6nHRwt9YPrgowBXhF7caT 0.122923
Claynosaurs 54n7eQ2N4mX2XCJCdVQ2NCqawAEtxibuAWYhm3auvLA 0.083121
Table 10. Portfolio for 99971.smap
Serie ID Weight
99971.smap 87nMGosNGvar8ynFEVdkW9fKenYPNuC5df7mou4yJenP 0.101152
99971.smap 5ZfZYvxx5Yycg6KcrVrFvYEDRjrE84Z8nh4xD8A3DEwGLo 0.082697
99971.smap Gfhe4o6SFCjEdH96PYqNZ3egNw517VCkTgEt5bxxULJ 0.040014
99971.smap BoLV4gG3ibokmIetJZcmzrtijJVSNvxrukcU2RP4kKr 0.13749
99971.smap 76KphRLzqp7gJDJYj4oLR2tvWA6oiQxJBNuqA2XaMM32 0.142183
99971.smap 3EW8XwacRyrUNJWMFF3dhNYz2xGfoTzYGsZ4i1abWR6mD8 0.058078
99971.smap 5sQHoU8gNJtdsypneDEGmtUwEwzLiPiSVsC37r3oTHuo 0.108522
99971.smap 4kuxRHVQ29p8zf vjWH6PGrDPDeFZ4WBPWhWGNz4aY3hS 0.123808
99971.smap BFAdmfVLP5DLkwC8bpim4kvsLkmJSgrzRuchJY9VbEn 0.098907
99971.smap BYxcioTS37me9sQwDGYN9GbCw3xiSaLEexRBDl1ks3nFV 0.107149
Table 11. Portfolio for BoDoggos
Serie ID Weight
BoDoggos 8Kawzhw97zRMU2EQBobeVwnP8TPxAr4W8BhynVhkwPFQ 0.074899
BoDoggos DhJmRxwmxdU37xY8vBuYkrswbuRUZUBqLw71uesD7x37E 0.128103
BoDoggos ClvehxPQYsiCpVhisyFB7KLgaPT7Bdm7iNjJnkpxRcKk 0.107006
BoDoggos 9NZjKEkxMo13g6TrMuAem5KrfnMXqAEUFVxmrrgFH5EUup 0.104183
BoDoggos 5ffFi15JoaAteC3C9KLRt57tuKGPdZdSo8cmAYx8Js4ip 0.11721
BoDoggos Gwjs8cQTUWbcVpbJUaYT9PHHnnRudtPVgyb8wRXUzSJTP 0.044183
BoDoggos 7YMuXWAtFUDx5JRGjpg6Pw6AeawCuaXNke5eNMB7jBCv 0.068985
BoDoggos ElbxPSxgaB3exJhanb7tsFSeiJqwMxPKNL618zcYhuD6 0.059639
BoDoggos AvYiFeBpEDsGQQQuKFD48qFGVq7SupyibxFK8B6cXFVB9P 0.161498
BoDoggos ExcrctS9eJs2P71D8d1qdKaJD9sFZDHUADjd3J2Hxlo 0.134295
Table 12. Portfolio for sharkx
Serie ID Weight
sharkx 8ZVd1F2nhtmTeJ8DiFU9UTcqoTEhrGeLihn7py6bmEZiK 0.104882
sharkx 8z2EkW87szi8FHnvaSdwmiJjiQyQk4Ed55x6TJk4sM7C4 0.006608
sharkx BuGSZqH5dtvJBR7bQ82PdQHZ2dT3HhVBPVQxmfsw5zNLp 0.117628
sharkx BA24a3dXkrWp2jFVnnv5S4VrLDsi38f7wgy1Ks8zjsyK 0.085837
sharkx FNBYQ9MtKh1x4AovUE9gGoWo8tQ7FRkz9raYhHuXaVzl 0.08783
sharkx D7PndXPu7PxGKT4V2zDKAgeq8zbhNJNWPr86GNJsh3qJw 0.111395
sharkx 9cphRzndeed4st8PzNvWtZQXP5eexecJ8562hXfc9tZVR4 0.103839
sharkx GRfKmsCAhCkT2GY9WXceDDox3Ja8XnNNxxWfBVzcC4DERM 0.129588
sharkx A8zP4866ycFmFCUxJsJm9FmxOfm5DwYHuMjpR6bmxT 0.088348
sharkx HwylAwyd6m3gQY7DKWeJJuF7ktGGom9qxNUriFWc416Q 0.164043
Table 13. Portfolio for Okay Bears
Serie ID Weight
Okay Bears 4Khd6FTpxDb2Efh32F9oadnPHg66H8pbpEgYSihvV7UB 0.121919
Okay Bears FsLB5wnMsJ2xYCZGqq4S9Enb8z6ZZZHLxdVWXLN1MEUR 0.110083
Okay Bears 3cKV3z9PaUQabkBmJx4JK86VGGuGoTC86eVGV7vJuwTTzo 0.106892
Okay Bears Bs2rB5KxsIqf4iGUjGZRrx70cMVueRcPcsDfEe7u6Rekz 0.088204
Okay Bears GCYFanRJjzbigUwkSo2JVafqtXXNxBUcCes8y362YJnR 0.07378
Okay Bears J7zf129gD9FzgGmoapFwqWQG9CXSKK6XiWpEFkFxU557 0.139859
Okay Bears 9ouE9Wff8EnGdbJwtX2f4xFnZdSCjpJnvNo2Eky90uXHpb 0.116198
Okay Bears DZrigBryMnShugJzgDGE8S54VadZr7Uieyk9FBnV7VHWRh 0.070201
Okay Bears 5Y8BYMsa7HyZ8F3912CLevNJPJxrya4SH17ZsgFQXVU55 0.090539
Okay Bears 5jNKE8TtNBvssrcrBvbAwpIFUrfBF56ZYiSpEkFxUv057 0.083105
Table 14. Portfolio for Gaimin Gladiators
Serie ID Weight
Gaimin Gladiators 3JWJ8y7LZz6nicpNxZr9air3KoCrCkJBoTExC3Kttrs 0.02785
Gaimin Gladiators 3kSdHd6cNDEC5jbSHLXpuUTr4sZLAPiQdsbqaJa7iqKHZ 0.178734
Gaimin Gladiators 7iMWzU63EYP4StHvpxP6XfWnsAPJwbhUVZmDw8SseTbZ 0.067891
Gaimin Gladiators AHq9JwQZYGCHVor8DEJTiJU7V1IfpQ57mQbb2vJdwG93 0.098778
Gaimin Gladiators 3KZj3nNRbpJtMTmkPoPWKrXkPGWn4D8HGcrHwiJ1pnM4N 0.108086
Gaimin Gladiators 87U3wEeSo2SvDTzE4gEkVaMAFygDZOfTbtBhdZrEc8S8PF 0.04587
Gaimin Gladiators C2S9UanArx8zxPWqLvgmDdF7t1xh3qrDg5vTFBCtXK287 0.161676
Gaimin Gladiators 7ZaNKLYVVVM8dZR8rTt8UYFUqKTa4HMSXxgzd2o62vzjc 0.024872
Gaimin Gladiators GDMcbQxn2hXQb5bgFIHEbQgbCjU2DMoJzAanNqcQH4u4y 0.166706
Gaimin Gladiators 5TEqinhyCyo1Lo2ocxnyW9ekCx2xM18rz37gLkkLi1RqUR 0.060821
Table 15. Portfolio for SMB Gen2
Serie ID Weight
SMB Gen2 7D8zmMEf6NXi7PKVTSrnAxmjp3i4aCmnFcVGoktGGgjrDV 0.105106
SMB Gen2 GBTz8ry2XKnvTshAA7ogtZfPPPBm3ZsLxtEDQSj9qgf2JB 0.075154
SMB Gen2 9GKzidBP99vaATtM7c8qmorMmnA8Tbncea8FJUkYSFBRc 0.133103
SMB Gen2 79DRTJ951qbiHn7FVqQmo46eAmxGGioeDeMcUeup5WXZ 0.056754
SMB Gen2 A7RzRyiMPppWX29rEJ7DXCjJppMAT1fKLsUEx8UiRUthH 0.110816
SMB Gen2 6S3iBdCaQ96CDobBqdWMQYv7p8T7c7y6ag2HtwpfKwQHA 0.078785
SMB Gen2 CC1dLYX3ezisa4HyXYjkFyvBS84SV5e5exoQomavJ9XW 0.097836
SMB Gen2 9Td2XnUirhccde495v79rWsoWzoJPJHDPCpkjcg3rD9ZRi 0.079573
SMB Gen2 GfW9dMsmwPdtGmMD9DK78zgeeWVevfKCVXtor9LMxdvG 0.099464
SMB Gen2 CgxgjEvHMRd51Lr37ig7ibITuRoctNKeQdb9dxf41Y747e 0.163408
Table 16. Portfolio for Peppermints
Serie ID Weight
Peppermints Fd4m9utB9QPRpmCVcnqpmJX2MbasiSAax95quXgTJRnU 0.153019
Peppermints G39RDisKusyLCa2TtxzeQFB5dUgD6hrBri8MqnaWy2PxX 0.056553
Peppermints C7Y5WT9pj7nmYbBxvggBDeWLox7ENiUWUB2JLsS9ZuwLsb 0.140747
Peppermints DWqKPGSvd2dIUmJkrYXTBCcbJAw4fFAwpunnFZjhNKUMP 0.111962
Peppermints 3uDZp3fh7KDRQPKLD5sGWm6bjw2kmV5sMG6Kv9kCPQjh 0.080909
Peppermints HFlvYs3ySZs1XL3fvEHCbc5Cyw3qGkbfStaTXocKYLHEM 0.103693
Peppermints 3E8e5vx8kv2RMXuUaKACYWNhttxrxcJ9qGsJrEtySovKe9 0.0578
Peppermints Es5oeodYClrmEWQe8enqBitKv2FA2jyJdueyipMGvYrL 0.102021
Peppermints 4Y7tQWGksWY7P45tjbW47nEjFYCAhpnQlc6Gi89T4NeL 0.06758
Peppermints 9ss3J6aA3Mm3PRtYMenGbwBNaeMPBnGcmTXotKndX5 0.117565
Table 17. Portfolio for Stylish Studs NFT
Serie ID Weight
Stylish Studs NFT F23AmcudzvQiNTqty4e2aGqKKBFoKVfLrr8xHEUqT7Q2P 0.109453
Stylish Studs NFT EoBP7V44Rkg1JLJg6qxiMd1MYJjhnN3PXsNuCYHunNq7BZy 0.10452
Stylish Studs NFT DEDMdMoUiHY1ipsqQWhzPNHkZvgB3LWDT2vrkm5CcgG7Ue 0.049525
Stylish Studs NFT HH5k6ggfoePjbxB4AWnqJYDkCF9tBdchP1QA18YbUiuc 0.123038
Stylish Studs NFT 8s7dV1EBJgcJ5uVsY4RTmzdEBEjkmM8JLQYtgaPR6S8 0.105151
Stylish Studs NFT HrQmb3macSxdvQv5WZqu3PTSJTaZ7RfMAFxb4bTLrtzf 0.09527
Stylish Studs NFT ETZevQZRLSon6Ee1b5vrAnVrD56n6P2hCK4qUstWq8A 0.133372
Stylish Studs NFT 9q55QoqtQxrxMWdFV1rLq9KnbY2mtrfWua2aPFFZMHeA 0.077666
Stylish Studs NFT 4aeC3Yh7SKRhnmqYaWHoPDEBcXBCKBHWrXDxpopAby61 0.121346
Stylish Studs NFT 9MVwP45t8CxnmZSPwAFdEzioUKCEUuNxRMdXkWViCwjb 0.08661
Table 18. Portfolio for ONE
Serie ID Weight
ONE GWunN4oyx49R4ZPtgM8Wve3aR9z46qtyb3k1itGtWJfD 0.095844
ONE 5aqhyfTZyPZaUEYDyo5set9U2ttMCM6Y2ZnDDFFUMgkv 0.03885
ONE 811FCFRdafQtQwgATCERjrighYQoE26Toyx7yFTZ25UQW 0.079332
ONE 2Hxgv9481PdmKVTELOblYHJVRF9Dc3vq15EHLZyb8xXf 0.139256
ONE ELjw3AmxEFGHwi7nu55em29vMAc4Mdw6vXcakJ7pYTM5 0.069998
ONE 99yVF8sA9DYM9FnaxU8kSAzdbn1DqghMUXpiYvccxcbcM 0.108488
ONE 95wvftv8MZxrECyqyFt4wSmHiefeePgjWKYu5p9mdLb 0.111751
ONE B1ep4gRBU2FY2SxyYZWfxfp8yJBLhgrnuNe8aFV42Zbt8 0.106138
ONE 2XbhNCyki4TMNmcm5N2v5KszwmJ8ALcVTFX3HAjoY1xy 0.128005
ONE 2EnkU2UH9j4YApcp9SkZcd6aXd9UTL9x2P8XQaWaUd 0.122338
Table 19. Portfolio for Kanpai Panda
Serie ID Weight
Kanpai Panda Bqs7YyFVFXCLUWj9Ukrbi5iGQVQ1bRJ9xyd6BfJY85Vqy 0.109178
Kanpai Panda DCYduWdgWFtEFlx9FQNwZnfp7cya4je19yDqSeKrZuRf 0.123128
Kanpai Panda Bl1hRGHRpDUd9MXyw89d98UrdC7jbwu3mTLWYk2icZbA 0.12257
Kanpai Panda J3dtPVka6yx9ZDGCeCju4vKR8qaoef6RLiY7ya9aX2YF 0.139829
Kanpai Panda GwjjMQ6ArjUEmPiZrPpGGCnpLb1j5s64cGj9J8xaTA4ZGLp 0.076419
Kanpai Panda 9thHb85XrUNoDZrASmxkDLWNYhaH3FssWLBwA3av4ERe7o 0.083098
Kanpai Panda E4Vgm3EVFLbcewcwdLzjaJeUJs9zoDExNq22LYewAcf97 0.069634
Kanpai Panda FUXbVsYG6aigZXxNfdsGT4xhhkGmhGkK1BszwZ49jFDal 0.093523
Kanpai Panda 46yewnycSaXi4QuBpm8Vqw9bFDpoZmiT2LPerLw8yhx 0.082198
Kanpai Panda 4xliKNGDnrxDh7FDYbuJ6tSW5YZQPKCFAy71babmVT7B 0.100423
Table 20. Portfolio for Quekz
Serie ID Weight
Quekz 6Lo59MeHj6esVPufQfsFGnVxKi7HomvJfuVhzFshcLSog 0.08562
Quekz AwSZTUMLn48kjbMbnZJZc61zf1j8XKF4GbN2Ukesra3K 0.112305
Quekz CuCDPoTM73JuT5gQF89bgrQGCNZaoncnhVWtWUiGnjS4 0.102783
Quekz AF7EBprNPHHFt8qMnTHuXdhzEz8ied4uP63quXBX7cfz 0.051148
Quekz CNym9SZ8xRefzBomHq3LM8M44cqmWBiebD3BbX96tgeo5 0.103646
Quekz 8tQd8kx82K6orf3VBe98KgNtgtXNW7qBRsLrvctzusQz 0.102167
Quekz HFBRuy7Hfni7Fn1yEkqGW12rMKSoPSHBTWZa6X4hWaqQ 0.119826
Quekz DU2RKyP9p315pvXRiXf3Y5zTNirRLK8AAC2hiwNaDL9Cp 0.117304
Quekz 4YVtsvmzN1Cu6okrvVPt9ccglLhgUEVTiufmDtTpxLVErt 0.114137
Quekz 2gyAUazag7Zc8iAYHQqofQNZYMT3o8jFH3eW7qEd6Eq 0.0901065
Table 21. Portfolio for Solcaino.io
Serie ID Weight
Solcaino.io 8ePKrdrV7HxPBM6Sv4B6oawy8Rd4v6Buao3kyabhMjJ1 0.108975
Solcaino.io 3Hraw6QooqEr9wcGMdHB6htGANDqJAVkRdRzlyuzZubh 0.123276
Solcaino.io 2WXyJuRqtRXgn2D3xWB5jjKvxqzubxegAysxWoKEiDLf 0.128963
Solcaino.io 4EhEdsKeT6T9TNnXYpZqonMdqotZS2hDiDuUJMRSi6wy 0.061094
Solcaino.io HRTFyPbPgPwqkMWv796MBuguZ1FxG62gPxSSAZUnVGn6M 0.122819
Solcaino.io GK9EVGiR74ktsRTSyUPJuZ02go3R34rgE1xATeElmvEgEp 0.094049
Solcaino.io 8RVNkDrKhRbpWDTA6vxLbxHL8CgC5ERMJEotaWWNSJLm 0.103723
Solcaino.io CAsFBol2SUzpSuz9pthLRzyQsHeuyusPb2NC25UHa5qfA 0.150778
Solcaino.io 5aWxV5v9zT0DxbTnTFu9MNDuFsHyJzeBRyTDn3xeAh 0.070626
Solcaino.io 7KvCs6uGHkgyaHXd3izne6Jzwd945muSYEijnEYuWMV 0.080697
Table 22. Portfolio for ZMB
Serie ID Weight
ZMB 5cLPKHNJ6L5jpPNwzVRT8KUMbTpgBLm6gbMJpYh5b2Nz 0.130018
ZMB HnQUQFPMbkKzb47qzKKQYsoxqb4B8zMkmu7TtgZtqoM9s 0.057949
ZMB 7SREzoqNWKM4sRDgHuejU6MgTqbUNTnhvv8sawS7YkqTje 0.10972
ZMB 4coZjiipdWPHA5fLSCSqP8yGLtpreuAaungUgEKg2rH26 0.076038
ZMB Hj5v29CAMaPz5EEiAoNFLPa7mPixZDPKmgowDbnuBwh1 0.086951
ZMB 2PYvdpbjZzAd7ZXCBmVmncuCMYTS8ny1pyieJTuDzLJw 0.125473
ZMB 3DCcf6fF7DEj16kVXTCadktvakNCNnzjVseTPt7wkKuCS 0.068714
ZMB 95YQPBI D9S2M3qxvwWnAyL4E6EcxxgsCVCwjoWFs9JAK 0.134505
ZMB 2qxQeteP7AqbnMpwoXRAAAXkFyZrPpmgmwBN9PdKzk9sa 0.122761
ZMB YeZ3xHv899sjkQhQ4Wfr3ZC37DvMk2wRwrV1sfvUEWZ 0.087872
Table 23. Portfolio for Stylish Stud
Serie ID Weight
Stylish Stud EMQmzVxDEtiwbbEwjLdndBQHwa7QNK8NviWZQwty26ag 0.065087
Stylish Stud FrtwvPxe2tgtltsWvRGMGQuARjyg1Cm4N78rgzfCF7sHa 0.168
Stylish Stud BTLEkiHAnpgWiUApaU9ti5GYCyGP6fvThmr8C8sFlsxv 0.129167
Stylish Stud 3cnYHKSKFiFjDdyBveiu7CGi6akege7sy6tqkZjzjUL7d 0.065365
Stylish Stud J3KBhKMGnQVL1F8eYUtCGrYRnSHpJ1Rh6jZxVazWQ9ns 0.057824
Stylish Stud D2WEMDtsVSKxtLe24ScAFzR72hZ6ZrXugPbhhuZeoWRnKi 0.108448
Stylish Stud GaJdfhVMLHFhgkVJCzxcRc87V2qeKeuFQhPrXZaPBoDSn 0.151372
Stylish Stud AdGXbWD3J4s4t7xnGsG6C3JYPR4ajyrP4wEMMDBFWzoY 0.123507
Stylish Stud 8Zx4vRNmABm16sBM522tLvnJPPGoFacsaQv78vZ4oe 0.016748
Stylish Stud EMhaxiuctner2Yd5x8voTbdSKHKCrk78xUbM7id19rVB 0.114482
Table 24. Portfolio for LIFINITY Flames
Serie ID Weight
LIFINITY Flames 32SWbwS4fmEb9MeJLFRek3C7ELACGrsmr2hmTJJ6fX7o 0.063232
LIFINITY Flames 8vFFZu7zZfntaWXdS19QWPLHzQQ6XdPB3VitmYpaW6jf g 0.125966
LIFINITY Flames CETJ2rX1xpjfmJ5jfAqqLiXcXtk8py5oyoWLffF1fAaU 0.088817
LIFINITY Flames 94Q6WinX57z51ovxcNTi4ye5ZyoMS7Eg7CjxxXP42vmC6 0.096272
LIFINITY Flames 7ijTPoHMfRL7mtwvWtUmaoCn4naGeSyF16cgov3di2i 0.091837
LIFINITY Flames A7ivXx2xATZJwJwJC8CincHd9FvWqF4vMXapPch11Ap 0.12131
LIFINITY Flames HvLzsYvekYyaJ4Tpv6qawnz2dpF1qgB2BuHxnj2js9CPcU 0.173131
LIFINITY Flames F4Ey15Uy5EgXVBAckw3SSSdzqVtdh5k3zMIHDUTapAhgW 0.069732
LIFINITY Flames AiyaGt75LWbwZ jZzdRDBX5rsCRPvCd8s1gTbmWzkCraVX 0.120512
LIFINITY Flames 7q1rRLMknLRNxZabWGCin41N8Yk7v6nCB3MdSSRQzKh62y 0.049192
Table 25. Portfolio for Galactic Gecko Space Garage
Serie ID Weight
Galactic Gecko Space Garage AGkdTcHFasnsvH4ofrcK5FXasmvghnl1ZSBHcKGj5BrDDW 0.094517
Galactic Gecko Space Garage 2XNg9ScN6vyMB7F5kdNZdwc6w7Fa3CD3Oeix5s5xnJbV4w 0.150622
Galactic Gecko Space Garage 2K6yCDMWUNLMvRB55S1XaJymKUpKqzivQyerFVoeWP6 0.122612
Galactic Gecko Space Garage B7aemXNu4qA3dudis7H9am3KG96m6m3CqkSClcTpTF3 0.120258
Galactic Gecko Space Garage 2TXdfDLyrkCtm7Di1hmf6xbR9DHYjggpJSia8ojr9QYbL 0.08115
Galactic Gecko Space Garage 39Be9GmugkMEwsAysdzmW3hmLVxnJ7BNHBnPfaBvk2jmv4 0.094866
Galactic Gecko Space Garage 5igAAxFmxwADkiMCzYD3yvPBxJYD5ELES1X3KBMMZsa4p 0.123284
Galactic Gecko Space Garage Henzzr8C2kdkMetBct51wMJQ5Z9u1krRB2u4lcDUJEgti 0.067053
Galactic Gecko Space Garage CZtembLM4J5jUWzYFpNWp2bHEqDml5Wzex2WwBleR 0.102909
Galactic Gecko Space Garage 8TzDOJDecyaxkWAeedM9ZLX7RXiF4EeqasXAjQmq6EQ 0.067017
Table 26. Portfolio for Degen Ape
Serie ID Weight
Degen Ape 99uc7fJh32vHwRwffi8Ncnrr3ByRZL3SQLvTfwyth73D 0.180677
Degen Ape BfrSvmwhrJb1rA6uFHYRWCU4P6Ad8mHLEpC6G2YvN1JR3 0.048066
Degen Ape 4v7v7YTGD6JXSRubJggtN26zxAeK1bmMoakWCJjy2ef 0.051087
Degen Ape PCjkYa9pH2HF26CJxgwc8EozadjuKoBgAv7r3bsrbfo 0.132624
Degen Ape FN4JE2zpR4Y2kc88SVQXcm22FryTfynWiaPtg9BxHq7H 0.108004
Degen Ape 8KYVXjxBV3fSnHg31xBYReu6ggz3Wzwr9jwUx4J2Abbn 0.035049
Degen Ape 9nEqQri6E7Jf7z7d9Ag9AqGMfkcE4UFGkYYtcP25UKAM 0.072588
Degen Ape 7FerQswjAYuhcadobW2jJ8RUWRb65cnWmm2X26w8hDi 0.13676
Degen Ape 4EvdRYmBVWMaAGcu6T3Tioq6hmfrZh8E8bYbWxBfS31n 0.107147
Degen Ape 62PNDR6e5RLk1RnZ7jHZSYW5vquBjuHlhqHgN85Gn3qR 0.127996
Table 27. Portfolio for BLOOD. IO PASS
Serie ID Weight
BLOOD.IO PASS 549sCo6tUpJz8h84gAGN776nAPE5nvaRAna6ksfTMJ3x 0.097675
BLOOD.IO PASS 7NFjK9c4XNknNxn7fwBVzLlkHCFsSTCz86mrjNeHRsQna 0.05709
BLOOD.IO PASS xmozPJKJZXlhvug72hPjK2EqXiDD5hQGPYc26WgZjAB 0.10367
BLOOD.IO PASS 9cHuTKYq1g1YigVDJQiauEzn8uPpGDSSgl1v9iAbiGz4Q 0.075953
BLOOD.IO PASS xDmdEZQShcDmwN6mZG9A2wQ42DKVSKFzWrmARubryeVqK 0.112238
BLOOD.IO PASS 7C7rV1ntUvTmhY8VvJzLj5KEsmsqcBFYkmE2O8v7Vmp1M 0.088471
BLOOD.IO PASS HkbAm18SdfrtyCZ8nD9tQBNbid3zpDZSNckE7n4zDRY7S 0.160822
BLOOD.IO PASS AQAMkzUu6GfKfKTEsLQoQWsJBFDeUE6ZGPhnNS12FRP8zc 0.095984
BLOOD.IO PASS JAVVYg9zr44fLhUEe6tAhPQzyhJ9PwK2CXiDx4qhfmj 0.097559
BLOOD.IO PASS 57GjCAtR5rJRspEQWdsT4cxJkUzSTLxpAPJqpfmpTwt 0.114539
Table 28. Portfolio for Alpha Gardener
Serie ID Weight
Alpha Gardener 9NXbDTNwdK8WgAophPT92FnwkCQAueVBuYhpe4qcbZoWu 0.065208
Alpha Gardener 3g8nJczivZbzB7jF4zuK9Lh2XKHvToS9RACBgN2jngTVZ 0.073699
Alpha Gardener Bkb6tocSoC52UXeEkHcMVCkhj8eP9Ucgf111v4xzXGYy 0.117214
Alpha Gardener 9xDV2wDFQThJrxvaQ1DtVcnooma3TueSxyC7rY3jJ7mNX 0.12907
Alpha Gardener EvksKQboACmPkpPsVYjVTuCsV7BhtEBzBKDRarri8BRQ 0.123471
Alpha Gardener 7WYvBbJuaNEMUeHGBzk37LSiyJB8X8zrosT5vOpcmx6wxd 0.093729
Alpha Gardener GqVdwuzuYgFDebJ2TccV6AGsCSepPFShpVRquRxpPy8Mw7 0.097566
Alpha Gardener HJi5k2pd7uVS8gweAwuTmpthAbnzcCZpK6i d5BnKDQwK 0.077987
Alpha Gardener 2V68ChrInBpU1 jvq jSqk iQJdmQc7C22o6ummmkPscBLcw 0.064803
Alpha Gardener 6vhNbZecGUL6HXZ9habQFW4f2JBKTgpt7xt8Z6KQmJDZ 0.157255
Table 29. Portfolio for Dappie Gang
Serie ID Weight
Dappie Gang AGpTHKh7jQ88RieRru3CX6B3MAnLrWbaBXbiXLEWa 0.033969
Dappie Gang Czboi98HQDqNEmgcuCZRRpg5iFVh4NeAXC4TLoiUo8sC 0.150486
Dappie Gang 4NxRZFg5xqqr7aF25VEMtwogTNByGfJ7hwXwd3NFwXyv 0.11618
Dappie Gang 3QQEmd3rDyHK4UDNwDmJhmMFfcvY4EcfBHAW4Bpxntml 0.093857
Dappie Gang 9bQtR4E82DJW4v4JWkx2WSc4jCbdQ5y1P1Cdljfe64YJ 0.099955
Dappie Gang zTFneSZ3DNAmnTRD3J43iV5iHq9Em7ky9R9Gbw2N51w 0.06051
Dappie Gang 6MUX81btYzXdDZLeh7yW3bS1FG659rNYzyShzx6RrPFXZd 0.103286
Dappie Gang GFXzggvG5ULusQaPZU n37dWAvpmLaJVPCSVQqt5KHCD6 0.130882
Dappie Gang E9E2zsLg3k1hWQGCfk88SBv7eAZrVzt2LGp9Uis2cAwgX 0.053654
Dappie Gang 47kmKRsBUNuH5VFeRcmew8taGNNcmGAm6vGLin3WxUuXi 0.12168
Table 30. Portfolio for Aurorian
Serie ID Weight
Aurorian 4cfN0b1bvpTBJVj262xmaKoaN8ChkqetdzVGTzsNu6ex 0.112044
Aurorian 2sEZextiyBMUNgXxsvedbztasgYmLRdd2jP72rWv4DoW 0.109016
Aurorian UKs64RK3PpdLFgTP6moCiqu5azvbL2wkBSbPptzzrTz 0.103812
Aurorian 2VABsSZzxPif4qjeSNNmovHnna33RWSNC19artKuS1CjJ 0.122614
Aurorian CSuAGuyWVJjroJnxgaqGC7CukBU42BttmYzbYVUs7LAMj 0.110863
Aurorian F1P6PjeA9cg5eiMAxEm8ECN43tDWkgGt4aHK2ATBZyBLF 0.091849
Aurorian Hd2aJeeinf3PDLbxRPYr7Giz3FLZsAa2nXKWHDg1PN5Z 0.080195
Aurorian 3LhvSNBziD24cWHDKgj nMJKvxOtxn2CrhuwZTYBosGMj 0.120744
Aurorian 9tQ4odLB1LmLPwzzxw3UxRfNUFCmSG9SMLWDKdvcS27W5 0.082134
Aurorian Fxdi4VFSQ9QLrAhyZBKvrMgdAeKgTncKvawsVccnKeCs 0.066729
Table 31. Portfolio for Ovo1
Serie ID Weight
Ovo1 8Z7rqyW1ysmMj6gRuFvJJC4eh4dtZegjaJYV18rfYfsC 0.151064
Ovo1 CPDeCW579LqTWKacVJGUCGLxonJMyeXuJrFF6ZRefYtsx 0.082613
Ovo1 5uFF5DF9MecSyYDHQQQTVt3esX4BxDkY6NGXC35GyCFJ 0.093908
Ovo1 3NV6rgZSDmha5CukqW51bBo5JdHNH8SauLNkm2tFuJhpx 0.071181
Ovo1 8TGEfMB5EGUE6HV4tGhCJoZW8vJRzBLztf1eWUP4WhWM 0.060748
Ovo1 A18kef6TT2XK8F89kSTMEhEnXY6SC9GxPswC6tyxxAq 0.044942
Ovo1 5mTFaoPBiSvCCNLACXJmDDpilrfnjbEZMT8Vv43mJ4LM 0.098761
Ovo1 FDZfastuaqVeaBvFvrnDmAoZfeEekdcqz28seasGMZcpy 0.154855
Ovo1 CkCYGWMqBTFvLD1WujrJV7kVNmNKvjd5UsV45JJnE 0.106808
Ovo1 6GntnwW54odvJN3HsaNalmLkn0XLhR7sq3RD6GGQNf3Gz 0.13512
Table 32. Portfolio for Beast
Serie ID Weight
Beast BNQ7gVRZuNZlJ97ekUrBGvpjvpnDwY1wLbCvhnfb9CJD 0.110164
Beast B2uBRGr3E6Ngkw75mEbppiMUE3qzMcY81MQYFi1Qtt8T 0.051987
Beast G3cDvmipcgLQWL8ichJWsdDDEgV2XpaGMz2sSK6Ump 0.061447
Beast Fcq2a5PdoPCT7gQAba2YyGWgH8gv46NGZYEpJE5qJ1dj 0.109936
Beast CdYWiLrqPoWMmrMBxETtruapE56SMVWNTJzq7AxnjM3G 0.118709
Beast BhkXpwkrjgbC8RGlRRoxslrxutNTL2ehpmEBU5SYDx7G 0.107245
Beast BUSUZ8cXBXPDUwNCzUSgDSgJFLHhW7RyWMJUuw55FHZax 0.171164
Beast 913DduqicDK6kW5yJMLpFj1i7YzP8mjlhWWAGJP3LPFpq7 0.085823
Beast TyaCEYcfuCM4wcxyYTb66uJjN3j5f7iQpAibfyQqncsMA 0.120421
Beast FyTSumanEb2nxvHeDnhCvOyJ3fnszhzGr8WWpR2wCg2i 0.063102
Table 33. Portfolio for CHADS
Serie ID Weight
CHADS ATgUWP5YZD6uW9XFqJTXsN2ZmNUrDzjpBT9XdaATxEoR 0.1486
CHADS 56fWQA24932Y21jeFVGCpXUi1Qo5xxy9LCpwnEJ6Rvox 0.10444
CHADS 4ggAkkp59je8tHkm9d4QxVh8fB6fwYCTHNNrDnn9uzdS9 0.117526
CHADS 2rdf7kkZK6ioyfwfyNc39agEnWrFHnXC5UfpcibG3rjLB 0.094137
CHADS 756g7kXvb3rEto4hH3sXe1HZEdTVbJaqcP4CZY7ggBPY 0.091999
CHADS E3Gq8ofgopobAnn8TqKo7bxnFB6shCTfEvSn69bTM2rP 0.07625
CHADS 8D9wrbHfcybnStd498YCZ6QY7a5i6ez5icBUantZAkWP 0.04402
CHADS DfbWUBJi91vuY23LwPERMJxrXSugt5e3DTAghmcnTjQN 0.096069
CHADS DKAorJ52vyEtCLJUDa1AzjSmWaPc7zEiicHGW9RGdxjb 0.146888
CHADS s16uuB6KAjHUoUX2rvuoYb8Y4urEstk7e9VDBTj6agQVT 0.080072
Table 34. Portfolio for Asset Dasha
Serie ID Weight
Asset Dasha FE68p8Ea8DxRwsiSe31atnG5KBqzxJERZsR247i4ACrm 0.045206
Asset Dasha BTosgHu2faqFL2SbUim4e7dhdBmRlurdEFHs49V2ifLf y 0.103462
Asset Dasha BdQMPmDz6DWmnQsyT6vJqeXY8KFkKH3wCVpeAuq4EVt 0.0849
Asset Dasha ASHE1BLWaXz4amdDFw44cNLX1SYJbUbE8wYiSKeJP4b 0.086867
Asset Dasha 6xjVoWkSFJ8YbnxBbCLNdJmAiR4poeQGumanJNPqxV5 0.11176
Asset Dasha BNttyKa5jiX49CjXNS5C8sPsaKU7dbfUKQKejV6GsQhQb 0.100946
Asset Dasha 9kTemokN2dEazeVn4taAC2QzME1SF6jNhND4si6PGroB 0.131874
Asset Dasha 9rhzuUZVpoKtJagFJvvBeL9KzgTu4xt7H75yLsk3v89k 0.121194
Asset Dasha 2mWHtB8pCdvozigLpRrR7tvnMUDufQyWcPBstzcZodg 0.101763
Asset Dasha 3rdlN6xtTfi7yqZFFPS8SmR2DyGCuWBQRKejM9TeBW 0.10628
Table 35. Portfolio for SMB Gen3
Serie ID Weight
SMB Gen3 25PsVjDupteheargbY5DVFR8yiH1sCbMksk5vi vmJ9qu 0.061138
SMB Gen3 G8G11AbgUWGN8kq3K3XtECpiBCnAugueruZHM6gfjVhe 0.077803
SMB Gen3 Bpr9zdbzzV1cwgpxEJb7aBXJlWWwJkxSAWmPBG7QuqE2 0.082595
SMB Gen3 9XMkaSXk1ebxs9ACp9u8WxkN7hhhozzyT4erCnvYBnqKH 0.128079
SMB Gen3 7HpQNNLsxNvWJc9RvaU29oY154nKEkvyv9EFrjFqN3yr 0.123526
SMB Gen3 4ckEyo8VstbMdqAmjorJaNCmxtpAEpAisEAnU9nE5yWi 0.085796
SMB Gen3 2PPicP4GKslY3C9Ryf4zA5DJnvco8XnCfePj5FscU6tT 0.100306
SMB Gen3 6q7YDMggh98Wdf4sMLPqChgCwBrH7uEvwPH1I bBR95UmP6 0.154229
SMB Gen3 7WqMa5Zn47y5TX98YYWeto4nLaMhRFHLC2yU1ENFW4Tq 0.07067
SMB Gen3 6JvSgn8Whl1EymRNCAQdEEqFsjMgU69N8Qg9Ezzw6SJP aH 0.109828
Table 36. Portfolio for Portals
Serie ID Weight
Portals 6qVAxFDswMqi28ExKaPJAsAbPkyrgxFcwPR6gR9qtngGDF 0.102339
Portals Evxm7bY5wgGh2vzaGgFuh8Attg2YlufnyecC5YNSzvnurUG 0.08097
Portals 67dt7j43JoEhrTcwnYa5GnhvYfzn8jEkakffu8nusXfJv 0.096435
Portals Fz7X12bewzbivXJbCZjz229QQH4Qdowt5jAMEyFZcBQm 0.113009
Portals 24jc6YR7TY6chmHME3wFv7mrU4wS9ZaJk5d6tHyVYgUH 0.105813
Portals Bods5gAga3BvmGVezzJjx6frbxDRzrnXvg55E9yxbSa2M 0.109874
Portals 7dLmHgqtWz2XP3kQTXY2MaRKVEVYnJmgYozTT8SoqU7W 0.15417
Portals Y4s9trWbCsQdWpmrzq5nYgm4mbC3yhecT869Z2Bfga 0.106054
Portals 8iSMJCBbfiTpZeRuEFSngUcKztKvPMC9c8BmAsDra9Qrt 0.053518
Portals 9h3GRBSfJo46CgPcMGAJBN4xToXB44B34JdI5n2cRzHNG 0.078185
Table 37. Portfolio for Blocksmith Labs
Serie ID Weight
Blocksmith Labs 8NFTVmahAHEmRdJBy1q9saGpJrAv2NQCPj9xktnWFiCj 0.05109
Blocksmith Labs 4tndcDLKpFcKk1qBdDlqH8Koxrrsg3sxgT5s6BD4DYEt 0.092024
Blocksmith Labs 9g4XUAQ9wBpEgJzyjwvRLihSbt1bF9dXinJScdCbE4s 0.073231
Blocksmith Labs 6UjULCzMBUZf8KaVJH7vb59i jtceH9pdTdeZdwK5gZ1n 0.122439
Blocksmith Labs FSYJzxFHddhxLvxYCAW8CkbUuBLx5PBFbAyi2VKLN55ye 0.10808
Blocksmith Labs EvCjD129EhsZfGHzSK3kmD95gsP0PQeEW7ZkNizelfQ 0.116293
Blocksmith Labs E8Eylscqoh6CNwRE9dfWbnHrQ1aU3p8buRfsP9pg6v4ae 0.108617
Blocksmith Labs 7PF2csNs4KmLSta2FYgnKcEr6XMnoahS6x3Xu6aKgjf1 0.122095
Blocksmith Labs 13zNBFmveVRiocs4cXBrD2olWZ10CMfEusecS4ims9Z2d 0.081388
Blocksmith Labs 5azJn2ECEiyRvGm1Ks3vc4zWDdfJaJKMaNR8kzNgUYv 0.124742
Table 38. Portfolio for Ticket
Serie ID Weight
Ticket BDj45c9yAFAsZQPF7VwJ5GNmCU5wEBkgHHAky4NP9Bw 0.111737
Ticket 3yuyuhdo8J9YhMnztRW5a jgtZVmroM7VkvPsyNLAvm5B3L 0.08276
Ticket 2XRsLKUfaPETuBJeSo5n8MUmLmHPzBYYwY6cDtbmDGFU3 0.120411
Ticket 5kDJ8qjpQwNaz4eMPwMXGHP182dcdBZvQfByQArFM1u9 0.114333
Ticket 7KVznzkoGjfnAnbgUXU87gEmr51SNwcYptkejb2cRzSx 0.090501
Ticket 4WS22coX3UNG LrNEcStokgxJxUdWefmJu92SU RuTaXb9 0.114754
Ticket 8mrFUDSSFSVRaeR4c9FTYkmAqgHiVBSU124mrUJ2ok54Rm 0.125037
Ticket 6SpZp3YsGp83YhPL8F jmm9KBuCbZinNinXwhnbgoNVR9 0.090018
Ticket Gbh6hVRunS6R1s8uYwjnKglBmnZLK7vzDxdvaJymZNj 0.078877
Ticket E8VEVaGHytUh6o2Cxo3Lnx1UcbBAX1pkdXUnMtJbshm 0.071573
Table 39. Portfolio for Clanynosaurz: Call of Saga
Serie ID Weight
Clanynosaurz: Call of Saga G6kaAepM6bocXC5aYzFBCLqwinBQaRg2V64CcmXYvqgCj 0.158797
Clanynosaurz: Call of Saga 2d3xnYXiqx9cdrYbZ5qYlDbko6NjBgePashnzEWykMCZMD 0.094572
Clanynosaurz: Call of Saga 40KB7Hxa4HF7xbpbrHxnPLMlRDLTmqMFuYDKgeKZLde 0.082894
Clanynosaurz: Call of Saga U1BfrR5tBKFPKsXnFtg3YJDJjEnvEsouIn2CM5BVEOde 0.120413
Clanynosaurz: Call of Saga 9QTmsTVE8s3gUjdYNNMwZ2sUjfJe5pMV357ozK5AnthKK 0.081728
Clanynosaurz: Call of Saga 8aCM986Lw5gbpxbToKxrAX7fFAv9avxFKQYtsBuCoCpJ 0.088636
Clanynosaurz: Call of Saga 7k3bcBtmL6clgxZTfidbBWYTVEnnxNa6fX12eCxCvAYk 0.122418
Clanynosaurz: Call of Saga 95amPSZKg4teTYaA2bMw43Pwpm5S5EwpdLF65jqp2C6P 0.113161
Clanynosaurz: Call of Saga 80qnHAXPpyzWpAWnQ6dViJAcLBhQTnrHn3aw6CbkXtJia 0.074642
Clanynosaurz: Call of Saga HZfTnm2YUdNRiTGrnWlVQWhQRqHVspYSkT87f2wFsR 0.08074
Table 40. Portfolio for Bozo Collective
Serie ID Weight
Bozo Collective FBrcQHrSBjJ1F3gYmqkczjx2bJc1fjaKMtsTc4HMVsz 0.072152
Bozo Collective 6LoHYLUA5mcBAsdoeqiNwWsVB2EBESq4JQkzLFNV9vCe 0.12444
Bozo Collective GS34iNEgBm5AX77NzUg7vQrpqYmBgawhAMUjQk3iJKkx 0.146667
Bozo Collective 2aRigguvINVM7k972F36WxSxxiQ2uib1vFWY6zpK14bo 0.115924
Bozo Collective C26tzgg8RCoBiVj7ZbfZj6wJ5JEfS7hZLzoFh6TwCJrG7A 0.141768
Bozo Collective Dkg4eW2xiZB668c0FaWSbXR4jS3dfdnUcaCS6GQkmYrc4 0.085664
Bozo Collective 3mYkfTglkjCsN6tnWzSbPbQQ9UrTZFMPA48rE8TXPrssr 0.048039
Bozo Collective AMkaZvQF6A4r6Ss2ynr1132Sg3GXcEkL5WSJ5hoFHf8g 0.130551
Bozo Collective 4BbCxZwwpU3HomUxJnfNDLkXrodnIAawJYiunDlEyFZ 0.060608
Bozo Collective 5jyucVyZn23rcP92exedbCD8EledmqG334ah9iYrQfw 0.074188
Table 41. Portfolio for Meekolony Pass
Serie ID Weight
Meekolony Pass HzKbheXESqMu94MPSVrktFQ6V7spSpiIIUQsVojLW5a3x 0.079024
Meekolony Pass 4M6hvDq5PpFLRiuLnT4EQ1SVuyuYX3YsWtW47aoGnGq4 0.086602
Meekolony Pass CFzqYkhX8Q4PgE3sp3VWLaFdSpaVscTuiA4KaZQJWAL 0.077212
Meekolony Pass 3Xzw2jUWS944JHGjexcCyZNoweaCDPZtuP j7ynYB i8psd 0.110042
Meekolony Pass ahRS629gWTrZ2DhqWexJDecxeJKVMoQYoLasS6tguGhZ 0.125243
Meekolony Pass 3Yh1T1hLUzPKleH9TLHMdeJ4XjMp7fSmnD9rBAi4Mzxj 0.08987
Meekolony Pass 2XjNARRg4AFFf8xSVgRwqR2uKuahmn9nhnaKvz2KUiBo 0.162573
Meekolony Pass CgVJR4n6vzmGvN97xbuaGcyS5ChErzYcjnracX3Sqksx 0.085186
Meekolony Pass Caw5MSnybESpLyMXpPDcx3nW4oyMNwiLysbNWvtJzLF9 0.079899
Meekolony Pass 97mgBBTLWjJffkmLAGbZRRqaxptpmlMntcrC97aUgvF2ux 0.104349
Table 42. Portfolio for Fred NFT
Serie ID Weight
Fred NFT 8zHy9RSVaJGZYJcHm6BoJDgFta6bvv9NdnwZY7EieGvg 0.106062
Fred NFT GhybD6Rq42s4aLPd4s4xpV9aXxfJoJvtV4AS2ZqvyRzSoYP 0.067892
Fred NFT 5N5PjssakmtREtbW1fnyoGF4YbRApDSdy4uNwcD8ZYLm 0.137533
Fred NFT 63uJnFi5j7nomYeh6ZChmzyhHLyBHOMnFCooMX9zRR4Y3 0.13922
Fred NFT EsvMarcoW4SSBk4jJXt7YhYd9R6MFdfdXYbhLpVYvve3HkgF 0.100033
Fred NFT H4QhssSPEMCh5r7o38Xmc9TUmgGL5tKqavKfyXxv4LDer 0.004196
Fred NFT 7rYXG8HRkc jd48mmr2pXmgdZXxoFeHikm7sJfbQJDBVe 0.131646
Fred NFT 36yuUZjfmv18Lq2qmnQnSgtE9CAoKhrTATRRV56rom71W 0.091758
Fred NFT 2g7W9m217j3Lat91GCJGRZvDqaw8gYUCUWuusuMeG32Pa 0.097866
Fred NFT 6RSKhd341aDPsh9phhfNu43T3kSRY8ujPiLQsSz6FkD 0.126775
Table 43. Portfolio for The Heist
Serie ID Weight
The Heist 48WT7ea5YdspkY9NJGgWUsdRHRGloRT8PNdrbfZ1YarjP 0.083681
The Heist Cp9H7ohMFHSuktoErY46sVqkQQ7NzSCaeADZYhHJ2KY4t5 0.106894
The Heist B9Me9BVie4NUDZFJGhmMn7qaGQAPkQzrSQXUWPlJcbG9c6 0.158856
The Heist 5TuNJr5JaEalSgluYdhbWz48hky8wpxrBFuJr3Qs1R4h 0.108851
The Heist 5f8npl4FaikHTYo26d8Dq1cwtfH2WTJz36DCmdYDCdcg3 0.0682
The Heist 58eYsF3xWneka6n8eusDcmyce8u5geyxQde3t3PPfk5A 0.075521
The Heist 6bFrEQSP4QrxSLuHwahdwc j75hzwWYZkmybAsmBvYxNJ 0.133831
The Heist CDWwM9pcwk76eTm5FzAvZyNhTfSahTwcWLJpgpssaamxC 0.078515
The Heist 85p3b17gCvJPVS4jCJq4ukNbxsaCdMARSvh7MTosSHXk 0.09983
The Heist EWkDrxSFO2KqQXPsaACynBQpAeVjKegtWCbwBgKC7mYhl iG 0.085821
Table 44. Portfolio for Degen Fat Cat
Serie ID Weight
Degen Fat Cat GMkadUhh8Jrb4Jnzvonr7SH53ecJ84ZtgoK8E2EkHdJJ 0.069619
Degen Fat Cat 796ANQaMGutbLUFmZjsMb6AH29PWS H4UMT27WM7dEs8r 0.090749
Degen Fat Cat AeDeqXUW8o j92bP2mmBpPqAzU5wwwivyoUwzgMFFEhDR 0.101542
Degen Fat Cat EVAyhbjge5kwDMDLdgPSSchDgwvSbrerFAvqBqLZEtJ2 0.153879
Degen Fat Cat DuzbXWRtacZB7jwKgtBKo3HREMjSdQpYuBgxuEDX5rJV 0.057552
Degen Fat Cat Dbxp8eL2bVU54zwtRSwK9gscZ4UDwMCmn76ix6BAys 0.105039
Degen Fat Cat B4Mf1EbqXiA8gYQMg2wo76Cqc8tbfELNp7vhbMKbzXm 0.108898
Degen Fat Cat 2pqVYgn64uYQT3YrVfgCiPHvTbEqWiluHzkoXKg4sxQDB 0.104243
Degen Fat Cat 5Px4XzMohmdxBn8xiFY3JZZLfA9K1RwJS3CPM9bzJbY 0.092671
Degen Fat Cat 9Eug6CdjpBVJXGjbRKNLLZJiqk4hlnCGynYjrXfqKazP 0.115807
Table 45. Portfolio for Honeyland
Serie ID Weight
Honeyland GEDFPuf71HPVExRkT1FNVyLN2BNc8wiXmkpNqKMSAMRY 0.105055
Honeyland 7gVD4CRGjzz9JvemGNrAYf7ZzuKM2eVxpPZJCpKWYSjqQ 0.085069
Honeyland x7DuSH6ELCNHvW8AZF2hJgUJCcwfyAA6t6SukyiJbcK 0.121546
Honeyland 3kERR3nBi8GScUa7jU5qF6ECBAKAYf5cXVekFuJyFdhe 0.132242
Honeyland Bu9jRDxrZp8kUSGq3N9dQPdUDnVN1JeoDWjqfh1BN3aX 0.102705
Honeyland 9dDaShykmiiRrnmppmws6E9FhmG6G4xdT3rvB2WJCfprV 0.068198
Honeyland hk63aPXJHsmF4YY8xHUMF8PWMWrPFN6wHdEjpuiD9nF 0.073581
Honeyland 5gSY5uXKK1szRaa5uGu8GuEgJEcPS47mibGvwjCPWrRUg 0.084763
Honeyland 9PZrn9BuFS8P54Pq52oaw744GuRa5gv8FaEDWGYERjbAk 0.092272
Honeyland 3lhtMnuPPsKAo2n9NWkas2EvpSjpfvF4NaYE9CFCKeGHb8CX6 0.13457
Table 46. Portfolio for Frog
Serie ID Weight
Frog FhbREJcnLNzzTvnMjCKtsqmpgtgdEaCP2HuWlexfuhzQ 0.127664
Frog 9A8ssJcthuBE3k8Ltq415oPUQC9b96SFAR3fcVVPkPYlW 0.020927
Frog 93TAMnXfV53Lj6dLBuioPoPPDBruHbxQQuDD9VCVZzc3R 0.098313
Frog DXXawpwGCNVVuPBy7eRef3zPK5gDTONRrnkGvECWDjtv 0.162892
Frog DppnVJaE1DLoeblc36BJNw5CFG2qwdjAGYJFCIAiFwz3w 0.15539
Frog 7Cv26pcCRFNHBw146mSCYhjbmqP7GsuKpg5NriL4WJXy 0.125189
Frog 699yuLYDqnPggf4x49Nvagy2TtVNwx6bJF jBjN6k5EMMw 0.043736
Frog 6H9hrunVJN9DWApUzfNx9WxBteTR73T5kufqQYe81AZv 0.012197
Frog 4bumKdc8vch9CMZ3f6J69TtQUQCFi1Nnw1IKX8NeSTmR 0.131742
Frog 91sBrPS8HgLrzm5HEp4EzmaUUmUpkZTh8pk5HiHrXb 0.121949
Table 47. Portfolio for STEPN
Serie ID Weight
STEPN 86ByhDLxghHrzV9rDLB5hKoc5jvC8n4fDb jo77HYTSEDC 0.091218
STEPN GKN1z5KvdNhNh492sJgNDA2er6o1CDptmpRrWCT7nR2gf 0.12483
STEPN HmvnTJ8q7sQ2CJy2nTQzqigfvSCexf6Bzrpqi8DCgsAe 0.144232
STEPN FgKEmaqrRGUCTA9z97JJF69bK4AkmC6481QZJ8biptCC5 0.085015
STEPN AoC7StryrEFaeomiK4AZ2ASZqFQ1ExuP JrrZuPU6p2Nz 0.091028
STEPN 4PEtLXZztTAZrzac3rdLAu8rJpDwHS9LYBcTFSq6DUDX 0.090991
STEPN 2b84fXd7aJp68RccgAKjaeo14JognSCVTpvXh7KPme5 0.060749
STEPN EuuvBVQT4MoSByDGcSt9HCM1smdny99Hx3c68g9YdGFz5 0.081377
STEPN 6c9GryitdcJeJEe9uv5NnZ11FyNFZA3NpQg61A75F4XA 0.110797
STEPN 3nWMMNLoa jphaUiEkbJ16LuFXqnsX4SWdWiFzbf54IHgdm 0.119762
Table 48. Portfolio for Cega Super Sonics
Serie ID Weight
Cega Super Sonics AM5CSavePtu393 ji8wQFxXYBWUt6icndwmxnKn8FsgTSD 0.079342
Cega Super Sonics 4DqbcVB8BzMrv4v4ag8eXrRx945iyMGezMtTjwksskoOn 0.128018
Cega Super Sonics 6eWca89SEi791mLRiHEYfc18o4KW4WKZZqg9JntEuWiP 0.092108
Cega Super Sonics ABL8CRZMP2mk185moZvkoV1kHgriPvooyKL3y9RSEXLX 0.172269
Cega Super Sonics B9X4DhCvH1hmWYFzFy2FJhtbjmvtzzrsJmr29tSUwBCMK 0.088943
Cega Super Sonics GPMka2V8PWQwZcgaPQKiz72TR4jXCjfJsHyonJgQYBZf2 0.088631
Cega Super Sonics Gaa47qyFkhD5LM8xRSgvzrxrjl8mk2KBq6RBYcn9Jt5HU 0.088943
Cega Super Sonics 6wiY5pEyiUsUrsrrh2DRYtn3yScKfehcAPRCTLnHmfNF2 0.092342
Cega Super Sonics 2nkibSROtMGdsTBn3aXHhrFGaYAYASBgqAYesGWbLnF 0.05188
Cega Super Sonics bbxEtSMeG26MGagh5Sh5caS9V7sVJ8vrN4V5gzesYeqS4ag 0.117037
Table 49. Portfolio for TYR
Serie ID Weight
TYR Bai8t8SPwExtJwndid35LbW3YNS9WJgk9KkTCAAhBotrU 0.093321
TYR 2TdjXZYzogtVHfFxYxucjNJ25xXGohfawtRaBaQFXcQF 0.076138
TYR 7BLNnALEoezUzs9c28Lqni1kU8T4RnH471wX81Wm6vJp 0.078318
TYR BDEulscFTuWQQ9z8VjmAKjFr2eyuvuvAvBXefsjtduToVYG 0.138707
TYR 6rnERRj3iviTpoaqFfjkbKDigxnSh7gr5zVTjuvwHR5mZ 0.115078
TYR Gf7EkNDXibWmqiJazKA5ZDGozmRPLGZEhoVao tuoEZ7i 0.088054
TYR pCLhwZnbSwrcCSXtDubCa87RbkrvXhW83yoQ1s9qCqQ 0.063497
TYR ExgG93CX2LgMSKnpBhoSboQm3yCjFj1wLVPLMWwzC2eL7JW 0.125263
TYR 51WMe9JVxhifa9oLZcS1M89aqhs12TaZikNxHk2KcoZ9 0.071445
TYR 7BhUj41aoyrqaQy36Gc2vEDFzqel9QemgG2d4kjgNFd6cb39 0.15018